### v1

# LetsTalk: Latent Diffusion Transformer for Talking Video Synthesis

Abstract
* Equal contribution. Correspondence to Ruibo Fu: ruibo.fu@nlpr.ia.ac.cn.
Portrait image animation using audio has rapidly advanced, enabling the creation of increasingly realistic and expressive animated faces. The challenges of this multimodality-guided video generation task involve fusing various modalities while ensuring consistency in timing and portrait. We further seek to produce vivid talking heads. To address these challenges, we present LetsTalk (LatEnt Diffusion TranSformer for Talking Video Synthesis), a diffusion transformer that incorporates modular temporal and spatial attention mechanisms to merge multimodality and enhance spatial-temporal consistency. To handle multimodal conditions, we first summarize three fusion schemes, ranging from shallow to deep fusion compactness, and thoroughly explore their impact and applicability. Then we propose a suitable solution according to the modality differences of image, audio, and video generation. For portrait, we utilize a deep fusion scheme (Symbiotic Fusion) to ensure portrait consistency. For audio, we implement a shallow fusion scheme (Direct Fusion) to achieve audio-animation alignment while preserving diversity. Our extensive experiments demonstrate that our approach generates temporally coherent and realistic videos with enhanced diversity and liveliness. The project page is available at Let’s Talk.
{strip}
[Uncaptioned image][Uncaptioned image]
Figure 1: We propose LetsTalk, a diffusion-based transformer for audio-driven portrait image animation. Left: Given a single reference image and audio, LetsTalks can produce a realistic and vivid video aligned with the input audio. Note that each column corresponds to the same audio. The results show that LetsTalk can drive consistent and reasonable mouth motions for the input audio. Right: Generation quality vs. inference time on the HDTF dataset, the circle area reflects the number of parameters of the method. Our LetsTalk achieves the best quality while also being highly efficient in inference, compared to current mainstream diffusion-based methods such as Hallo and AniPortrait. In addition, our base version (LetsTalk-B) achieves performance similar to Hallo with only 8 $\times$ fewer parameters.
1 Introduction
Animating a portrait image with speech, also known as talking head generation, is a challenging task that aims to generate a speaking portrait from a single image with audio. Naturally driving the reference portrait according to the given audio shows great potential in video gaming, film, digital marketing, etc. However, producing reasonable animation faces the following main challenges: consistency in timing and portrait, audio-animation alignment, and diversity.
To overcome these challenges, several methods [52, 45] adopt a 3D morphable model (3DMM) [4, 26, 3] to represent the reference portrait and use audio signals to drive the 3DMM. They feed the 3D signals (e.g. coefficients, poses) from the driven 3DMM into a neural renderer to generate a talking head. While achieving great consistency in timing and portrait, they tend to produce dull results, such as unblinking eyes and fixed hair, which damage the diversity of animations. In contrast to these 3D-aware methods, some other methods [7, 32, 55, 10, 54] mainly focus on capturing the correspondence between the given audio and related 2D movements, like facial muscle and lip motions, directly aligning animation with audio. They implicitly learn the correspondence between audio and animation in the latent space, animating the portrait image by feeding forward the audio. However, these methods still struggle to disentangle the components representing the static and dynamic aspects of the face, even hindering the temporal consistency and diversity of videos.
Recently, diffusion models [18, 13, 34] have emerged as a promising technique to generate realistic 2D images [34, 51] and have even achieved remarkable results in the video generation [29, 28]. Notably, AnimateAnyone [21] incorporates motion modules and ReferenceNet into diffusion models to generate realistic and pose-controllable human animation, which verifies that the end-to-end diffusion models can seamlessly capture the correspondence between the condition signal and human animation. Inspired by it, recent methods [47, 9, 11] replace the pose signals in AnimateAnyone with audio signals and achieve talking head generation. Leveraging diffusion models, these methods excel at generating various unconscious actions, thereby enhancing diversity. While achieving promising results, these methods adopt U-Net as the backbone and lack the flexibility for better fusing multimodality. In addition, although the U-Net backbone works well for image generation, it introduces considerable payloads and struggles with scalability for video generation. Therefore, it is necessary to propose a transformer framework to solve this task and reasonably fuse multimodality, ensuring better portrait consistency and producing more vivid characters.
In this work, we present LetsTalk, a latent diffusion transformer to solve this video generation-like task via the diffusion paradigm. In LetsTalk, we first employ a variational autoencoder (VAE) to reduce this task into a generation task within low-resolution latent space. To enhance the spatial-temporal consistency, we equip each transformer block with temporal and spatial attention modules to capture intra-frame and cross-frame coherence. Regarding multimodal guidance, we define it in this task as the fusion of video generation with reference image and audio sequences. To better exploit multimodality to guide video generation, we summarize three schemes from shallow to deep according to the compactness of the fusion: 1) Direct Fusion, which uses a fixed modality to guide all features in the backbone through a cross-attention mechanism; 2) Siamese Fusion, like ReferenceNet, which feeds modality into a backbone-like network and uses the corresponding features to guide the features in the backbone; 3) Symbiotic Fusion, which concatenates modality with the input at the beginning and achieves guidance only through backbone’s inherent self-attention mechanisms. After summarizing these fusion schemes, we explore the impact of adopting different schemes for the given reference image and audio sequence. To better guide video generation, we use Symbiotic Fusion for the reference image whose modality is close to the video while strongly ensuring the portrait consistency of the generated video. Given the audio, we ask the animation to be aligned with it, generating plausible animation, especially lip motions, and further pursuing diversity (e.g. blinking, head moving, and expressions). Considering the audio modality differs from video, we adopt the Direct Fusion that ensures the audio-animation alignments and avoids ambiguity caused by significant modality differences in the fusion. Additionally, we find that unconscious portrait actions, closely related to diversity, are primarily generated during the denoising of diffusion models, and the shallow fusion scheme effectively preserves the diversity. The empirical results indicate that our innovative framework realizes realistic and high-fidelity talking-head synthesis, advancing the frontiers of realism and expressiveness in audio-driven video generation. To the best of our knowledge, our method is the first work to introduce the diffusion transformer for audio-driven portrait animation.
We summarize the contributions of this work as follows:
  • We present LetsTalk that pioneers the utilization of transformers in diffusion-based, audio-driven portrait animation by introducing modular spatial-temporal attention mechanisms;
  • We summarize and deeply analyze various multimodal fusion schemes, and employ Symbiotic Fusion and Direct Fusion to fuse the input reference portrait and audio respectively, reasonably guiding video generation via multimodality;
  • Our experiments on standard benchmarks demonstrate the superiority of our methods to other approaches in driving portrait animation via audio.
2 Related Work
2.1 Diffusion Models for Video Generation
Diffusion models have demonstrated exceptional potential across various domains, particularly in image [18, 13] and video generation [17, 21, 20, 57]. A notable example is Stable Diffusion [34], which utilizes a UNet architecture to iteratively produce high-quality images from text prompts after extensive training on large-scale text-image datasets [37]. This approach has been extended to more complex tasks like video generation [17, 21], where models such as Video Diffusion Models (VDM)[20] apply the diffusion process with space-time factorized training, and ImagenVideo[19] enhances VDM with cascaded diffusion models for high-resolution outputs. Further advances, including VideoComposer [43], AnimateDiff [17], and VideoCrafter [6], focus on synthesizing videos by conditioning the diffusion process on images and blending image latents. A significant direction in the evolution of diffusion models is the incorporation of patch sequences of images and transformers, exemplified by the Diffusion Transformer (DiT) architecture [31], which facilitates scaling to larger datasets and model parameters. Recent works, such as Latte [29] and VDT [28], have extended the DiT architecture for video generation by incorporating temporal mechanisms, achieving strong results by training diffusion models from scratch. Building on this, our study also adopts an end-to-end approach for high-fidelity, temporally consistent video generation.
2.2 Audio-driven Portrait Image Animation
In recent years, the field of portrait image animation has advanced significantly, especially in generating realistic and expressive talking head animations from static images paired with audio inputs. Early efforts primarily focused on improving lip synchronization, with models like LipSyncExpert [32] achieving high accuracy for specific identities. Subsequently, methods such as ATVGnet [7] and MakeItTalk [55] introduced facial landmarks and disentangled representations to improve video frame generation. The integration of 3D information in models like SadTalker [52], Genface [50], and DiffTalk [38] further improved head movements and expressions. A significant advancement came with the introduction of diffusion models, including Diffused Heads [39], DreamTalk [30], and VividTalk [40], which pushed the field forward by producing high-fidelity, natural talking head videos across diverse identities, focusing on improved lip synchronization and varied speaking styles. Recent innovations like VASA-1 [48], EMO [41] and AniPortrait [45] have focused on creating emotionally expressive talking faces, capturing a wide range of facial nuances and movements. AniTalker [27] introduced a universal motion model for realistic facial movements with minimal labeled data, enabling dynamic avatars. LivePortrait [16] extends the implicit-keypoint framework, balancing efficiency and controllability. Most recently, Hallo [47, 11] and X-Portrait [46] have adopted the latent diffusion framework pioneered in AnimateAnyone [21] or ControlNet [51], adding more control over elements such as lip movements, expressions, and poses.
2.3 Conditionally Guided Diffusion Model
The impressive generative capabilities of diffusion models have made the integration of control information during generation a crucial area of research. Text-guided control methods focus on refining prompts, manipulating CLIP features, and modifying cross-attention mechanisms [33, 1, 5, 49, 24]. ControlNet [51] enhances large pretrained text-to-image diffusion models by incorporating spatially localized, task-specific conditioning. In video generation, AnimateAnyone [21] introduces the ReferenceNet to ensure the consistency of fine-grained appearance details from reference images. Gen-1 [15] integrates structural guidance through depth maps and cross-attention, while MCDiff [8], LaMD [22], and MotionCtrl [44] focus on motion-guided video generation, enhancing the realism of human actions. For talking video synthesis, previous approaches primarily relied on image diffusion models to generate dynamic portraits frame by frame, starting from reference portraits [38, 39]. Recent methods often incorporate ReferenceNet for processing reference portraits and cross-attention mechanisms for handling audio information [41, 47]. While control over both images and videos in latent diffusion models has seen significant advancements, diffusion transformers remain underdeveloped in this area, particularly when it comes to handling multiple modalities (e.g. image, video, audio). To address this gap, we explore various diffusion conditioning strategies tailored to the transformer architecture, aiming to achieve better control and enhanced generation quality.
(a) Overview of LetsTalk
(b) Transformer Block
Figure 2: The overview of our method (a) and the illustration of our designed transformer block (b). For better illustration, we omit the timestep encoder and Layer Norm in (b). LetsTalk integrates transformer blocks equipped with both temporal and spatial attention modules, designed to capture intra-frame spatial details and establish temporal correspondence across time steps. After obtaining portrait and audio embeddings, Symbiotic Fusion is used for fusing the portrait embedding and Direct Fusion is for fusing the audio embedding. Notably, we repeat the portrait embedding along the frame axis to make it have the same shape as the noise embedding.
3 Method
We present LetsTalk, a diffusion transformer for portrait image animation driven by speech, as shown in Fig. 2. In this section, we first briefly introduce diffusion models and the formulation of our task in Sec. 3.1. Then, we delve into the details of our transformer backbone in Sec. 3.2. Finally, we summarize multimodal fusion schemes and present our adopted strategy for fusing the input reference portrait and audio in Sec. 3.3.
3.1 Preliminaries
Diffusion Formulation The diffusion process incrementally corrupts data $x_{0}$ with Gaussian noise: $q(x_{t}|x_{0})=\mathcal{N}(x_{t};\sqrt{\bar{\alpha}_{t}}x_{0},(1-\bar{\alpha}_{t})\mathbf{I})$, where $\bar{\alpha}_{t}$ are predetermined hyperparameters. By reparameterizing, we obtain: $x_{t}=\sqrt{\bar{\alpha}_{t}}x_{0}+\sqrt{1-\bar{\alpha}_{t}}\epsilon_{t}$, where $\epsilon\sim\mathcal{N}(0,1)$, forming a Markov chain over $T$ steps.
Diffusion models are trained to learn the reverse process, modeled as: $p_{\theta}(x_{t-1}|x_{t})=\mathcal{N}(\mu_{\theta}(x_{t}),\Sigma_{\theta}(x_{t}))$, with neural networks parameterizing $p_{\theta}$. This is optimized via the variational lower bound of $\log p(x_{0})$:
$\begin{split}\vspace{-0.3cm}\mathcal{L}(\theta)=&\sum_{t}\mathcal{D}_{\text{KL}}(q^{*}(x_{t-1}|x_{t},x_{0})\,\|\,p_{\theta}(x_{t-1}|x_{t}))\\ &-p(x_{0}|x_{1}).\end{split}$ (1)
With both $q$ and $p_{\theta}$ Gaussian, $D_{\text{KL}}$ is computed from their means and covariances. Reparameterizing $\mu_{\theta}$ as a noise-prediction network $\epsilon_{\theta}$ enables training via mean-squared error between $\epsilon_{\theta}(x)$ and sampled noise $\epsilon_{t}$:
$\mathcal{L}_{\text{simple}}(\theta)=\|\epsilon_{\theta}(x)-\epsilon(x)\|_{2}^{2}.$ (2)
To incorporate a learnable reverse covariance $\Sigma_{\theta}$, we minimize the full $D_{\text{KL}}$ term, following Nichol and Dhariwal: optimizing $\theta$ with $\mathcal{L}_{\text{simple}}$ and $\Sigma_{\theta}$ with the full $\mathcal{L}$.
Diffusion Transformer To better handle the multi-modality fusion in the talking head task, we are motivated to extend the Diffusion Transformer (DiT) in our framework. DiT builds upon the Vision Transformer (ViT) architecture, which processes sequences of image patches. This sequential patch-based processing mechanism natively supports the fusion of multi-modality sufficiently and efficiently. Different from ViT, DiT incorporates a variational autoencoder (VAE) [25] in the framework, thus runs in the latent space, rather than pixel space. Such design enables DiT to handle the expensive high-resolution generation task, which brings more obvious benefits in the video generation task.
Since we follow DiT and introduce a pretrained VAE in our framework, our goal is reduced from generating a video $\bm{V}\in\mathbb{R}^{F\times\hat{H}\times\hat{W}\times 3}$ to generating a latent feature $\bm{F}\in\mathbb{R}^{F\times H\times W\times C}$, where $F$ denotes the number of video frames333 The encoder in VAE downsamples the image by a factor of $8$, and $\hat{H}=8\times H,\hat{W}=8\times W$ We first separate the latent feature $\bm{F}$ into non-overlapping square patches along the spatial dimensions (i.e., $H$ and $W$) and obtain a patch sequence of length $F\times P$. For each patch, we flatten it and use an MLP to project it into the latent space as $\bm{Z}\in\mathbb{R}^{F\times P\times d}$, where $d$ is the dimension of the latent space. Then we add spatial and temporal positional embeddings [42] to the latent feature and make it sensitive to spatial and temporal information. Finally, we feed the processed latent feature into transformers and perform denoising.
(a) Direct Fusion
(b) Siamese Fusion
(c) Symbiotic Fusion
Figure 3: Illustration of three multimodal fusion schemes, our transformer backbone is formed by the left-side blocks. (a) Direct Fusion. Directly feeding condition into each block’s cross-attention module; (b) Siamese Fusion. Maintaining a similar transformer and feeding the condition into it, extracting the corresponding features to guide the features in the backbone; (c) Symbiotic Fusion. Concatenating modality with the input at the beginning, then feeding it into the backbone, achieving fusion via the inherent self-attention mechanisms.
3.2 Efficient Spatial-Temporal-Aware Transformer Backbone
Given a sequence of latent features $\bm{Z}\in\mathbb{R}^{F\times P\times d}$, we feed it into a transformer and predict the noise to be removed. For video generation, we extend the backbone in DiT that only consists of spatial transformer blocks. It is straightforward to flatten the $\bm{Z}$ into $F\times P$ tokens with shape $(F\times P,d)$ and repeatedly conduct a self-attention mechanism to achieve spatiotemporal fusion. However, this approach results in a huge payload for computing the attention matrix with shape $(F\times P,F\times P)$, as it grows with the square of the increase in resolution or frames.
To address this problem, we propose to achieve spatiotemporal fusion in an “interleaved fusion” manner, implemented by alternately stacking spatial and temporal blocks. Considering that $F$ refers to the temporal dimension (i.e. the number of frames) and $P$ refers to the spatial dimension (i.e. the number of patches in each frame), to feed the feature into the spatial block, we maintain the feature with shape $(F,P,d)$ and calculate the attention matrix with shape $(F,P,P)$ to capture spatial information exclusively among patches in the same frame. Similarly, to feed the feature into the temporal block, we maintain the feature with shape $(P,F,d)$ and calculate the attention matrix with shape $(P,F,F)$ to capture temporal information among frames in the same patch. Compared with the straightforward scheme that produces an attention matrix with $F^{2}\times P^{2}$ parameters, our scheme significantly reduces the amount to $F\times P^{2}+P\times F^{2}$. Our experiments verify the efficacy of our backbone extension in terms of image quality and temporal consistency.
3.3 Multimodal Fusion Schemes for Portrait and Speech Audio
After defining the transformer backbone, the next step is to propose a fusion scheme that exploits the input reference image and audio to achieve the multimodality-guided video generation task. Given the condition modality, we summarize the existing fusion solutions into the following three schemes based on the transformer paradigm:
  • Direct Fusion (Fig. 3(a)) directly feeds the condition modality into each block of the backbone and fuses it with the inner features via cross-attention.
  • Siamese Fusion (Fig. 3(b)) feeds the condition modality into another similar transformer, extracts features from each block, and fuses them with the features in the corresponding block of the backbone;
  • Symbiotic Fusion (Fig. 3(c)) concatenates condition modality with inputs before feeding it into the backbone and achieving seamless fusion via inherent self-attention.
After defining these schemes with fusion compactness ranging from shallow to deep, we aim to adopt reasonable solutions for input modalities (i.e. the reference image of portrait and audio).
Given the reference image, we require that the character of the generated video is consistent with the reference, i.e. using the reference image to strongly guide video generation. Moreover, Since the image modality is close to the video, Symbiotic Fusion is appropriate. Specifically, we first encode the reference image and obtain the portrait feature $\bm{P}\in\mathbb{R}^{P\times d}$ in the latent space. Then we temporally repeat the portrait feature $\bm{P}$ with $F$ (i.e. frames) times and concatenate it with the latent feature $\bm{F}$ along the spatial dimension [^4] To concatenate latent feature $\bm{F}\in\mathbb{R}^{F\times P\times d}$ with the portrait feature $\bm{P}\in\mathbb{R}^{P\times d}$, we extend and repeat the portrait feature along the temporal dimension as $\tilde{\bm{P}}\in\mathbb{R}^{F\times P\times d}$, the shape of concatenated latent feature is $[F,2P,d]$. and feed them into the transformer backbone. In contrast, we observe that Direct Fusion attempts to use shallow portrait feature $\bm{P}$ to guide the learning of all blocks, leading to misaligning the features in the deeper blocks. While Siamese Fusion addresses this misalignment and achieves comparable results to Symbiotic Fusion, it needs to maintain a siamese transformer similar to the backbone, which introduces additional model parameters and complicates the learning task.
Datasets Raw Filtered
#IDs #Hours #IDs #Hours
HDTF 411 15.75 356 13.95
CelebV-HQ 35,666 68 22,423 41.27
Train - - 20,501 49.69
Eval - - 2,278 5.522
Figure 4: The dataset splits for HDTF and CelebV-HQ, including raw/filtered data and training/evaluation distributions.
As for the audio, we require it to reasonably drive the animation, especially the lip motion [52]. However, certain unconscious actions (e.g. blinking and head movements) are almost independent of the audio yet serve to diversity. Therefore, we aim for the audio to have a minimal correlation with these animations. Further considering that the modality of audio is quite different from the video, using Siamese Fusion and Symbiotic Fusion often causes ambiguity in the backbone learning, damaging the diversity. In a nutshell, Direct Fusion is concise yet effective for fusing the input audio. In addition, for the audio representation, to accommodate the tokenization mechanism of the transformer, we follow [47] and extract richer semantic information from different layers of the wav2vec encoder [36]. A sliding window mechanism is employed to capture long-term sequential audio information. Through nonlinear mapping, we obtain audio tokens enriched with semantic content, which are aligned with latent features in the same feature space. Temporal-level audio representations are then synchronized with video sequences, resulting in the semantic audio representation. The experiments in Sec. 4.4 verify the efficacy of our mechanism in feature fusion.
4 Experiment
4.1 Experimental Setups
Datasets We primarily conduct comprehensive experiments on two public datasets: HDTF [53] and CelebV-HQ [56]. These datasets include videos that consist of individuals of various ages, ethnicities, and genders, captured in half-body or close-up shots against varied indoor and outdoor backgrounds. To ensure high-quality training data, we adopt the preprocess scheme used in [47]. We first exclude videos with scene changes, significant camera movements, excessive facial motion, and side-facing shots. Then we refine the remaining videos and retain only single-person speaking videos with strong audio-lip correspondence. After refinement, we obtain 356 clips with 13.95 hours from the HDTF, and 22,423 clips with 41.27 hours from the CelebV-HQ dataset. During training, we extract 16-frame video clips from these datasets using a specific sampling interval, each frame is resized to $256\times 256$ resolution.
Evaluation Metrics The evaluation metrics utilized in the portrait image animation approach include Fréchet Inception Distance (FID), Fréchet Video Distance (FVD), Structural Similarity Index Measure (SSIM) and Expression-FID (E-FID). Specifically, FID and FVD measure the similarity between generated images and real data, with lower values indicating better performance and thus more realistic outputs [23]. SSIM indices measure the structural similarity between ground truth videos and generated videos. E-FID leverages Inception features to assess image authenticity, providing a more precise fidelity measure. It extracts facial expression parameters via a reconstruction method [12] and computes FID to quantify the gap between generated and ground truth expressions.
Baseline In experiments, we compare our method with the publicly available implementations of DreamTalk (34M) [30], SadTalker (138M) [52], AniPortrait (2.8B) [45], and Hallo (2.4B) [47]. The evaluation is performed on the HDTF and CelebV-HQ Dataset, utilizing a training and testing split where $90\%$ of the identity data was allocated for training purposes, as shown in Fig. 4.
Figure 5: The qualitative comparisons with other cutting-edge methods on the HDTF dataset. Our method achieves better audio-animation alignment (e.g. lip motions) and produces expressive results.
Figure 6: The qualitative comparisons with the existing portrait image animation approaches on the CelebV-HQ dataset. Our method achieves better portrait consistency.
Method FID $\downarrow$ FVD $\downarrow$ SSIM $\uparrow$ E-FID $\downarrow$
DreamTalk 105.434 529.510 0.406 3.466
SadTalker 102.371 362.634 0.414 2.168
AniPortrait 78.284 331.117 0.623 2.903
Hallo 45.754 285.988 0.698 2.235
Ours (Base) 46.240 289.385 0.666 2.587
Ours (Large) 37.134 219.119 0.632 1.962
Table 1: The quantitative comparisons with the existed portrait image animation approaches on the HDTF dataset.
Method FID $\downarrow$ FVD $\downarrow$ SSIM $\uparrow$ E-FID $\downarrow$
DreamTalk 64.297 293.905 0.448 5.915
SadTalker 42.017 77.802 0.469 3.372
AniPortrait 33.408 107.236 0.578 3.722
Hallo 8.504 49.431 0.580 3.280
Ours (Base) 9.585 36.159 0.598 3.349
Ours (Large) 7.610 33.175 0.584 3.038
Table 2: The quantitative comparisons with the existed portrait image animation approaches on the CelebV-HQ dataset.
Implementation details Our approach differs from previous methods [41, 47, 45] that require a two-stage training process. We adopt an end-to-end training strategy. Thanks to the scalability of transformers, we set up models with two sizes: the Base model, which is suitable for simple fine-tuning, and the Large model, which demonstrates better generalization ability. Following common practices in generative modeling works [31, 2], an exponential moving average (EMA) weight with a decay rate of $0.9999$ is maintained throughout training, and all reported results are directly derived from the EMA model. Additionally, we utilize a pre-trained VAE from Stable Diffusion 1.4, which remains frozen during training. Please refer to Sec. 7 for more implementation details.
4.2 Quantitative Results
The comprehensive quantitative comparisons of HDTF and CelebV datasets are listed in Tab. 1 and Tab. 2, respectively. The results show that our method significantly outperforms other state-of-the-art methods in almost all metrics, achieving the best portrait consistency and audio-animation alignment. Notably, our approach achieves the lowest FID (37.134) and FVD (219.119) on the HDTF dataset, as well as the lowest FID (7.610) and FVD (33.175) on the CelebV-HQ dataset. These results show that our method generates both realistic and temporally consistent talking-head videos. For a fair comparison, each method generates the first video clip for every validation video, which is then resized to the same resolution for evaluation.
4.3 Qualitative Results
We have visualized the results of different methods on the HDTF dataset (Fig. 9) and the CelebV-HQ dataset (Fig. 10). From the visualization results, it is evident that our method performs exceptionally well on both datasets. In comparison to the visual cues most strongly associated with audio (such as facial muscle and lip motions), our results closely align with the standard Ground Truth. Benefitting from our Symbiotic Fusion, our generated videos maintain strict portrait consistency. Furthermore, by employing Direct Fusion of audio, the characters exhibit strong expressiveness and diversity in our generated videos while achieving audio-animation alignment. More results can be found in Sec. 8.
4.4 Ablation Study
In this section, we analyze the impact of different fusion schemes as illustrated in Fig. 3 on guidance using reference portrait and audio. Specifically, for analyzing Siamese Fusion, we have to design Siamese transformers for both reference portrait and audio to produce guidance for each block of the backbone. For the reference portrait, we leverage a DiT model pre-trained on ImageNet. For the audio, we define a transformer whose blocks are equipped with two-layer MLPs. The parameters of these transformers are simultaneously updated during training. We conduct experiments on the HDTF dataset using the LetsTalk-B model to validate the impact of different fusion schemes.
Analysis on the Reference Portrait Fusion Portrait consistency is crucial to the portrait image animation task and highly relevant to the portrait fusion scheme used. We evaluate the synthesis quality of portrait fusion using three schemes, and the quantitative results are shown in Tab. 3. Symbiotic Fusion almost surpasses the other two schemes, realizing better portrait consistency and video quality. Besides, while Siamese Fusion produces comparable results, it introduces additional parameters (i.e. siamese transformer) and complicates the learning task.
Analysis on the Driven Audio Fusion The way we fuse audio affects both audio-animation alignment and diversity. Audio-animation alignment refers to specific visual cues [52] (e.g. facial muscle and lip motions), whereas the diversity is related to the speech-agnostic motions (e.g. blinking and neck moving) that often appear as unconscious actions. We observe that diversity mainly comes from the transformer backbone during denoising. Thus, we aim to fuse the audio to achieve alignment while preserving the backbone’s inherent diversity. Quantitative results in Tab. 4 indicate that Direct Fusion outperforms the other schemes in establishing the correspondence between audio and animation. To validate the impact of different fusion schemes on diversity, we serve LetsTalk as an unconditional video generator [^5] Filling audio embeddings with 00 for fusion. and compare the diversity of character actions in the generated videos. The results in Fig. 7 demonstrate that the deeper the fusion compactness, the more static the generated video. The tightest Symbiotic Fusion generates a portrait with almost no motions, Siamese Fusion produces slight motions, while the least tight Direct Fusion we adopt results in natural motions.
Figure 7: Portrait animation results without audio guidance for three fusion schemes. Siamese and Symbiotic fusion schemes with high fusion compactness limit the diversity of video generation, while Direct fusion preserves this diversity.
Reference Portrait FID $\downarrow$ FVD $\downarrow$ SSIM $\uparrow$ E-FID $\downarrow$
Direct Fusion 52.999 389.01 0.497 3.149
Siamese Fusion 47.779 286.692 0.651 2.597
Symbiotic Fusion 46.240 289.385 0.666 2.587
Table 3: Experimental results of three multimodal fusion schemes, guided by reference portraits, on the HDTF validation dataset, employing the Letstalk-B model.
Driven Audio FID $\downarrow$ FVD $\downarrow$ SSIM $\uparrow$ E-FID $\downarrow$
Symbiotic Fusion 51.880 403.628 0.471 3.398
Siamese Fusion 48.041 318.637 0.585 2.706
Direct Fusion 46.240 289.385 0.666 2.587
Table 4: Experimental results of three multimodal fusion schemes, driven by audio inputs, on the HDTF validation dataset, employing the Letstalk-B model.
5 Conclusion
We present LetsTalk, a novel multimodality-guided diffusion transformer to achieve superior audio-driven portrait animation results using audio input. In LetsTalk, we introduce modular spatial-temporal attention mechanisms into our transformer backbone to enhance temporal consistency. We summarize three fusion schemes to handle multimodality and deeply analyze their impact. We then adopt the optimal fusion schemes for input reference portrait and audio, respectively. Our extensive experiments demonstrate the effectiveness of LetsTalk in achieving state-of-the-art portrait animation results with a good balance between performance and payload.
Supplementary Material
6 Network Architecture
6.1 Backbone
In LetsTalk, we borrow the Variational AutoEncoder (VAE) [25] from Stable Diffusion [34]. Given the reference portrait image and noisy video, we first encode them using the encoder in VAE. Then we patchify and flatten them along the spatial dimension. Finally, we use an MLP to project them into the latent space. The embeddings in the latent space will be fed into our transformer backbone. Specifically, our LetsTalk model is constructed using a series of N transformer blocks, each with a hidden dimension of D and H multi-head attention heads. Inspired by ViT and DiT, we explore two configurations of LetsTalk with varying numbers of parameters, while maintaining a fixed patch size of 2, as detailed in  Tab. 5. Via the encoder in VAE, the downsampling factor is $8$, the input reference portrait image $\bm{I}_{\text{p}}\in\mathbb{R}^{256\times 256\times 3}$ and the noisy video $\bm{V}\in\mathbb{R}^{16\times 256\times 256\times 3}$ are encoded as a portrait feature $\bm{P}\in\mathbb{R}^{32\times 32\times 4}$ and a video feature $\bm{F}\in\mathbb{R}^{16\times 32\times 32\times 4}$. Next, the two features are partitioned according to $2\times 2$ patch size and flattened as $\bm{P}\in\mathbb{R}^{256\times 16}$ and $\bm{F}\in\mathbb{R}^{16\times 256\times 16}$. Take our base model (i.e. LetsTalk-Base) as an example, we use an MLP to project both features into the latent space as $\bm{P}\in\mathbb{R}^{256\times 768}$ and $\bm{F}\in\mathbb{R}^{16\times 256\times 768}$.
In the spatial-temporal-aware block, we integrate the block details into our pipeline to model the relationship between temporal and spatial attention. Inspired by  [28], we introduce a fully connected layer between the temporal and spatial attention mechanisms. This layer applies a linear transformation to the output of the temporal attention, enhancing the learning of temporal features and ensuring consistent dimensionality for subsequent processing.
Model Patch size Layer num N Hidden size D Paramaters
LetsTalk-B 2 12 768 193M
LetsTalk-L 2 24 1024 663M
Table 5: Details of LetsTalk models. We follow ViT and DiT model configurations for different model sizes.
6.2 Audio Encoder
We resample the audio signal to a 16 KHz sample rate. For audio feature extraction, we use wav2vec [36], employing embeddings from the last 12 layers to capture a richer range of semantic information across different levels of the audio representation. To ensure frame-by-frame alignment between audio and animation, we match the audio features with corresponding frames of the video clips. Additionally, we use a sliding window (multiple audio frames) for each video frame to maintain audio continuity and preserve more accurate audio information. To extract audio representations suitable for transformer tokenization, we designed an audio projection module. This module first flattens the extracted audio features and then selects a fixed number of context tokens (32 in our experiments). Simultaneously, the audio representations are mapped to the same dimensional feature space as the latent noise.
6.3 Siamese Transformer
The Siamese Fusion scheme has been adopted in AnimateAnyone [21], EMO [41], Hallo [47], and so on. The Siamese Fusion scheme aims to update a backbone-like network simultaneously and extract the features to fuse it into the backbone block-wise. Since these methods follow Stable Diffusion and use a U-Net to predict the noise for denoising, they adopt a similar U-Net as the siamese network. Considering LetsTalk as a Diffusion Transformer, we introduce a ViT [14] pre-trained on ImageNet to achieve the Siamese Fusion scheme for the reference portrait image. For the input audio, we propose a transformer whose blocks consist of a two-layer MLP to implement the Siamese Fusion scheme.
7 Additional Implementation Details
7.1 Diffusion
We leverage a pre-trained variational autoencoder model from Stable Diffusion 1.4, with all our diffusion models operating in the latent space of this VAE. After sampling a latent from the diffusion model, we decode it to pixel space using the VAE decoder. We adopt the hyperparameters from ADM [13], specifically employing a linear variance schedule with $t_{\text{max}}=1000$, ranging from $1\times 10^{-4}$ to $2\times 10^{-2}$. Additionally, we use ADM’s parameterization of the covariance $\Sigma_{\theta}$ and their approach for embedding input timesteps.
7.2 Long-Duration Generation
Figure 8: The illustration of the long-duration generation process.
Unlike the previous frame-by-frame generation approach [38, 39], our end-to-end diffusion transformer generates a video clip in a single sampling process. In diffusion-based video generation tasks, ensuring proper temporal alignment is critical for maintaining coherence and consistency across generated video sequences. Additionally, due to the enhanced realism and dynamism of the talking videos we generate, direct transitions between consecutive clips are often not feasible. To address this, we follow [47] and pick the last $N$ frames from the last generated clip as the initial frames of the current clip ($N=2$ in our experiments). The initial frames are solely involved in guiding the temporal attention, ensuring continuity between consecutive clips during sampling. Fig. 8 illustrates the long-duration generation process.
7.3 Training
We initialize LetsTalk from a pre-trained Latte [29] model on the FaceForensics dataset [35]. Following the training strategy outlined in Latte, we adopt gradient clipping to stabilize training and prevent gradient explosion. For training, we use the AdamW optimizer with a constant learning rate of $1\times 10^{-4}$ and no weight decay. The batch size ranges from 10 to 16, and the only data augmentation applied is horizontal flips. Each model is trained for 300k to 500k iterations. Training is conducted on both RTX 3090 and A800 GPUs, using automatic mixed precision to optimize both performance and memory usage.
7.4 Inference
During the inference stage, the network takes a single reference portrait and driving audio as input, generating a video sequence that animates the reference image based on the corresponding audio. The length and resolution of the generated video clip during inference need to remain consistent with those used during training, which in our experiments are $16$ frames and a resolution of $256\times 256$. We use DDPM for the sampling strategy, with a total of $250$ sampling steps.
8 Qualitative Comparisons
In this section, we report comprehensive qualitative comparisons of LetsTalk with other cutting-edge methods and demonstrate more applications of LetsTalk. We show more results on the HDTF [53] and Celeb-HQ [56] benchmarks ( Fig. 9 and Fig. 10).
We show the portrait animation results of LetsTalk driven by English (Fig. 11), Chinese (Fig. 12), and singing audios (Fig. 13). We further use the powerful AI assistant (Doubao) to generate portraits for comparisons (Fig. 14). Note that LetsTalk is only trained on the HDTF and Celeb-HQ datasets, which only contain daily conversations of real persons.
Figure 9: The qualitative comparisons with other cutting-edge methods on the HDTF dataset.
Figure 10: The qualitative comparisons with other cutting-edge methods on the Celeb-HQ dataset.
9 Visualization
We provide video visualizations in our attachment. Notably, the model used to generate the video is only trained on HDTF [53] and CelebV-HQ [56], which only include English conversation scenarios of real portraits. In addition, we use Chinese audio and singing audio to drive real portraits. Furthermore, LetsTalk can drive AI-generated portraits [^6] We use Doubao to generate AI-generated portraits in the attached video.. The visualization results demonstrate that LetsTalk has great generalization and achieves remarkable results outside the dataset.
Figure 11: English audio-driven results of LetsTalk. Note that each column corresponds to the same audio.
Figure 12: Chinese audio-driven results of LetsTalk. Note that each column corresponds to the same audio.
Figure 13: Singing audio-driven results of LetsTalk. Note that each column corresponds to the same audio.
Figure 14: Results of LetsTalk driving AI-generated portraits using English audio. Note that each column corresponds to the same audio.
10 Limitation and Future Work
Despite the advancements proposed in this study regarding portrait image animation, several limitations necessitate further exploration and consideration. (1) Resource and Dataset Limitations: Unlike recent works such as Hallo [47] and EMO [41], which use diverse real-world datasets, our study is limited by the availability of publicly accessible datasets. These datasets are relatively small and may not be as robust, thus limiting the generalizability of LetsTalk. Additionally, the computational resources available were insufficient to fully exploit larger models, restricting us to smaller-scale models with fewer parameters. Future research could benefit from larger datasets and scaling up model sizes to explore the scalability of LetsTalk. (2) Long-Duration Animation: A key limitation in generating long-duration animations with our end-to-end framework is the challenge of maintaining smooth transitions between video clips. Due to the high realism and dynamism of the generated clips, seamless continuity is difficult to achieve. While the method in  Fig. 8 mitigates this by using motion frames, the discrepancy between real frames used during training and synthetic frames used during inference leads to cumulative errors. As the video length increases, these errors accumulate, resulting in a degradation of quality in the later clips. Future work will focus on minimizing this error propagation, possibly through improved temporal modeling or alternative approaches to frame synthesis during inference. (3) High-Resolution: Although Latent Diffusion Models (LDM) [34] leverage VAE to enhance the resolution of synthesized images, high resolution remains a significant challenge for video tasks. Recent works in portrait animation, such as [41, 45, 47], have circumvented this issue through a two-stage training process. However, we hope that future research will explore more effective solutions for generating high-resolution outputs in end-to-end models like LetsTalk.