### v2

# Efficient Long-duration Talking Video Synthesis with Linear Diffusion Transformer under Multimodal Guidance

Abstract.
Portrait image animation using audio has rapidly advanced, but challenges remain in efficiently fusing multimodal inputs while ensuring temporal and portrait consistency with minimal computational cost. To address this, we present LetsTalk, a LinEar diffusion TranSformer for Talking video synthesis. LetsTalk incorporates a deep compression autoencoder to obtain efficient latent representations, and a spatio-temporal-aware transformer with efficient linear attention to effectively fuse multimodal information and enhance spatio-temporal consistency. We systematically explore and summarize three fusion schemes, ranging from shallow to deep fusion. We thoroughly analyze their characteristics, applicability, and trade-offs, thereby bridging critical gaps in multimodal conditional guidance. Based on modality differences of image, audio, and video generation, we adopt deep (Symbiotic Fusion) for portrait to ensure consistency, and shallow (Direct Fusion) for audio to align animation with speech while preserving motion diversity. To maintain temporal consistency in long-duration video generation, we propose a memory bank mechanism that preserves inter-clip dependencies, effectively preventing degradation across extended sequences. Furthermore, we develop a noise-regularized training strategy that explicitly compensates for DDPM sampling artifacts, significantly improving the model’s robustness in continuous generation scenarios. Our extensive experiments demonstrate that our approach achieves state-of-the-art generation quality, producing temporally coherent and realistic videos with enhanced diversity and liveliness, while maintaining remarkable efficiency through its optimized model design with 8$\times$ fewer parameters.
Talking Head Generation, Multimodal Guidance, Long-duration Generation, Linear Diffusion Transformer, Deep Compression Autoencoder
ccs: Computing methodologies Procedural animationccs: Computing methodologies Computer vision tasksccs: Computing methodologies Neural networks
Figure 1. We introduce LetsTalk, a diffusion-based transformer for audio-driven portrait animation. Given a reference image and audio, LetsTalk generates realistic videos with synchronized mouth motions. As shown in the Left figure, each column corresponds to the same audio, demonstrating consistent and accurate lip movements. The Right figure compares generation quality and inference time on the HDTF dataset, where circle size represents model parameters. LetsTalk achieves superior quality and efficiency, outperforming methods like Hallo and AniPortrait. Notably, our base version (LetsTalk-B) matches Hallo’s performance with 8$\times$ fewer parameters.
[…]
Enjoying the baseball game from the third-base seats. Ichiro Suzuki preparing to bat.
1. Introduction
Animating a portrait image with speech, also known as talking head generation, is a challenging task that aims to generate a speaking portrait from a single image with audio. Naturally driving the reference portrait according to the given audio shows great potential in video gaming, film, digital marketing, etc.. Previous works (Zhang et al., 2023a; Wei et al., 2024) use a 3D morphable model (3DMM) (Blanz and Vetter, 1999; Li et al., 2017; Bao et al., 2021) to represent reference portrait and drive them with audio signals. These approaches input driven 3DMM parameters (e.g., coefficients, poses) into neural renderers, achieving temporal and portrait consistency yet lacking output diversity, often showing static features like unblinking eyes and rigid hair. In contrast, audio-to-2D facial motion methods (Chen et al., 2019; Prajwal et al., 2020; Zhou et al., 2020; Cheng et al., 2022; Zhou et al., 2021) directly align audio with facial dynamics via latent space learning. However, their limited disentanglement of static and dynamic components compromises both temporal consistency and animation diversity.
Recently, diffusion models (Ho et al., 2020; Dhariwal and Nichol, 2021; Rombach et al., 2022) have emerged as a powerful paradigm for generating realistic 2D images (Rombach et al., 2022; Zhang et al., 2023b) and have demonstrated remarkable success in video generation (Ma et al., 2024; Lu et al., 2023). Notably, AnimateAnyone (Hu, 2024) introduces motion modules and ReferenceNet into diffusion models to produce realistic and pose-controllable human animations, showcasing the capability of end-to-end diffusion models to effectively capture the correspondence between conditional signals and human motion. Building on this, recent methods (Xu et al., 2024b; Chen et al., 2024; Cui et al., 2024) have extended this approach by replacing pose signals with audio signals for talking head generation. These methods leverage diffusion models to generate diverse unconscious actions, significantly enhancing the richness of generated content.
Despite these advancements, several critical challenges remain unaddressed. First, current approaches suffer from suboptimal efficiency in audio-driven animation, primarily due to: (1) reliance on auxiliary spatial signals (e.g., predefined movement regions) as a crutch for limited audio controllability, which sacrifices motion diversity and naturalness; (2) redundant parameter overhead from fine-tuning generic text-to-image/video models, despite their misalignment with animation-specific requirements; and (3) inefficient architectures, exemplified by low-compression-rate autoencoders and vanilla attention mechanisms with high computational complexity. Second, despite the focus on conditional video generation, there is a lack of systematic exploration into the intrinsic modeling principles for effectively integrating multimodal conditions, such as reference portraits and driving audio. Lastly, while existing methods attempt to achieve long-duration portrait animation through sequential frame synthesis, they fail to address the progressive accumulation of residual noise in generated frames. This phenomenon inevitably introduces cumulative artifacts (e.g., facial distortions or flickering) that amplify over extended sequences, ultimately degrading output fidelity.
In this work, we introduce LetsTalk, a linear diffusion transformer for efficient generation of realistic, long-duration talking videos under multimodal conditions. Built upon a diffusion transformer as the video generation backbone, LetsTalk replaces conventional text-to-image architectures with an efficient design tailored for audio-driven animation. Key to our approach is a deep compression autoencoder that compresses input frames 32×, drastically reducing the number of latent tokens while preserving critical visual details. Each transformer block incorporates dedicated temporal and spatial modules, where linear attention (Xie et al., 2024a) ensures efficient modeling of intra-frame and cross-frame coherence without sacrificing computational efficiency.
Regarding multimodal guidance, we define it in this task as the fusion of video generation with reference portrait and driving audio. To better exploit multimodality to guide video generation, we first summarize three schemes from shallow to deep according to the compactness of the fusion: 1) Direct Fusion, 2) Siamese Fusion, and 3) Symbiotic Fusion. We analyze these fusion schemes in terms of integration, conditional state, and parameter sharing, systematically summarizing their adaptability to different modalities, and further investigate the underlying factors influencing such adaptability—thereby bridging theoretical gaps in conditional guidance modeling. To better guide video generation effectively, we use Symbiotic Fusion for the reference image, ensuring strong portrait consistency due to its visual similarity to the video. For audio, we employ Direct Fusion to align animations (e.g., lip motion, expressions) with the audio while maintaining diversity in movements (e.g., blinking, head motion).
To enhance long-duration video generation, we incorporate a memory bank mechanism to maintain contextual continuity across sequential video sequences. Furthermore, we propose a novel noise-regularized training strategy and a patch drop augmentation to align the memory bank between training and inference. By explicitly simulating the residual noise accumulation that occurs during inference (where errors propagate as the video extends), our approach learns to compensate for such artifacts during training for robust long-sequence generation.
We summarize the contributions of this work as follows:
  • We present LetsTalk, a linear diffusion transformer for audio-driven portrait animation, which employs efficient linear attention for spatiotemporal modeling and integrates a deep compression autoencoder to achieve 32 $\times$ latent space reduction while preserving visual fidelity;
  • We systematically analyze three multimodal fusion schemes, characterize their adaptation properties, and employ Symbiotic Fusion and Direct Fusion to fuse the input reference portrait and audio respectively, reasonably guiding video generation via multimodality guidance;
  • We enable long-duration video generation via a memory bank framework with noise-regularized training, aligning memory states between training and inference to mitigate cumulative artifacts in extended sequences;
  • Our experiments demonstrate our method’s superiority in generating realistic, temporally coherent videos with enhanced diversity, achieving an optimal balance between computational efficiency and output quality.
2. Related Work
2.1. Diffusion Models for Video Generation
Diffusion models have demonstrated exceptional potential across various domains, particularly in image (Ho et al., 2020; Dhariwal and Nichol, 2021) and video generation (Guo et al., 2023; Hu, 2024; Ho et al., 2022b; Zhu et al., 2024). A notable example is Stable Diffusion (Rombach et al., 2022), which utilizes a UNet architecture to iteratively produce high-quality images from text prompts after extensive training on large-scale text-image datasets (Schuhmann et al., 2022). This approach has been extended to more complex tasks like video generation (Guo et al., 2023; Hu, 2024), where models such as Video Diffusion Models (VDM)(Ho et al., 2022b) apply the diffusion process with space-time factorized training, and ImagenVideo(Ho et al., 2022a) enhances VDM with cascaded diffusion models for high-resolution outputs. Further advances, including VideoComposer (Wang et al., 2024b), AnimateDiff (Guo et al., 2023), and VideoCrafter (Chen et al., 2023b), focus on synthesizing videos by conditioning the diffusion process on images and blending image latents. A significant direction in the evolution of diffusion models is the incorporation of patch sequences of images and transformers, exemplified by the Diffusion Transformer (DiT) architecture (Peebles and Xie, 2023), which facilitates scaling to larger datasets and model parameters. Recent works, such as Latte (Ma et al., 2024) and VDT (Lu et al., 2023), have extended the DiT architecture for video generation by incorporating temporal mechanisms, achieving strong results by training diffusion models from scratch. Building on this, our study also adopts an end-to-end approach for high-fidelity, temporally consistent video generation.
2.2. Audio-driven Portrait Image Animation
In recent years, the field of portrait image animation has advanced significantly, especially in generating realistic and expressive talking head animations from static images paired with audio inputs. Early efforts primarily focused on improving lip synchronization, with models like LipSyncExpert (Prajwal et al., 2020) achieving high accuracy for specific identities. Subsequently, methods such as ATVGnet (Chen et al., 2019) and MakeItTalk (Zhou et al., 2020) introduced facial landmarks and disentangled representations to improve video frame generation. The integration of 3D information in models like SadTalker (Zhang et al., 2023a), Genface (Ye et al., 2023), and DiffTalk (Shen et al., 2023) further improved head movements and expressions. A significant advancement came with the introduction of diffusion models, including Diffused Heads (Stypułkowski et al., 2024), DreamTalk (Ma et al., 2023), and VividTalk (Sun et al., 2023), which pushed the field forward by producing high-fidelity, natural talking head videos across diverse identities, focusing on improved lip synchronization and varied speaking styles. Recent innovations like VASA-1 (Xu et al., 2024a), EMO (Tian et al., 2024) and AniPortrait (Wei et al., 2024) have focused on creating emotionally expressive talking faces, capturing a wide range of facial nuances and movements. AniTalker (Liu et al., 2024) introduced a universal motion model for realistic facial movements with minimal labeled data, enabling dynamic avatars. LivePortrait (Guo et al., 2024) extends the implicit-keypoint framework, balancing efficiency and controllability. Most recently, Hallo (Xu et al., 2024b; Cui et al., 2024), EchoMimic (Chen et al., 2024) and X-Portrait (Xie et al., 2024b) have adopted the latent diffusion framework pioneered in AnimateAnyone (Hu, 2024), adding more control over elements such as lip movements, expressions, and poses.
2.3. Conditionally Guided Diffusion Model
The impressive generative capabilities of diffusion models have made the integration of control information during generation a crucial area of research. Text-guided control methods focus on refining prompts, manipulating CLIP features, and modifying cross-attention mechanisms (Ramesh et al., 2022; Avrahami et al., 2022; Brooks et al., 2023; Yang et al., 2024; Khachatryan et al., 2023). ControlNet (Zhang et al., 2023b) enhances large pretrained text-to-image diffusion models by incorporating spatially localized, task-specific conditioning. In video generation, AnimateAnyone (Hu, 2024) introduces the ReferenceNet to ensure the consistency of fine-grained appearance details from reference images. Gen-1 (Esser et al., 2023) integrates structural guidance through depth maps and cross-attention, while MCDiff (Chen et al., 2023a), LaMD (Hu et al., 2023), and MotionCtrl (Wang et al., 2024a) focus on motion-guided video generation, enhancing the realism of human actions. For talking video synthesis, previous approaches primarily relied on image diffusion models to generate dynamic portraits frame by frame, starting from reference portraits (Shen et al., 2023; Stypułkowski et al., 2024). Recent methods often incorporate ReferenceNet for processing reference portraits and cross-attention mechanisms for handling audio information (Tian et al., 2024; Xu et al., 2024b). While control over both images and videos in latent diffusion models has seen significant advancements, the theoretical understanding of various conditioning strategies remains limited, particularly in handling multiple modalities (e.g., image, video, audio). To address this gap, we investigate conditioning strategies in diffusion transformers, aiming to improve control and generation quality across modalities.
Figure 2. The illustration of our method. LetsTalk integrates a deep compression autoencoder that efficiently reduces spatial redundancy while preserving temporal features, and transformer blocks equipped with dual temporal and spatial attention modules, enabling the capture of intra-frame spatial details while establishing temporal correspondences across frames. After obtaining portrait and audio embeddings, Symbiotic Fusion is used for fusing the portrait embedding and Direct Fusion is for fusing the audio embedding. Notably, we repeat the portrait embedding along the frame axis to make it have the same shape as the noise embedding. To support long-duration video generation, we introduce a memory bank module to preserve temporal consistency across video sequences. Additionally, we devise a novel training strategy to to align memory bank between training and inference. By explicitly simulating the residual noise during inference, our approach learns to compensate for such artifacts, ensuring robust long-duration generation.
3. Method
We present LetsTalk, a diffusion transformer for audio-driven portrait animation, as illustrated in Fig. 2. In this section, we first detail our efficient transformer architecture in Sec. 3.1. Next, we outline multimodal fusion schemes in Sec. 3.2 and describe our strategy for integrating reference portraits with audio in Sec. 3.3. Finally, we introduce our noise-regularized training approach for long-duration video generation Sec. 3.4.
3.1. Efficient Spatial-Temporal Transformer with Compression Autoencoder
Deep Compression Autoencoder: Building upon latent diffusion models, we employ a deep compression autoencoder (Xie et al., 2024a) that downsamples images by a factor of 32 (vs. 8 in conventional VAEs), while maintaining reconstruction quality comparable to SDXL’s autoencoder through advanced architectural designs. This design is motivated by two key insights: (1) high-resolution video frames contain substantial redundant information that can be safely compressed, and (2) for audio-driven generation, separating compression from denoising enables specialized processing - the autoencoder preserves crucial temporal features (e.g., lip-sync dynamics) while the diffusion model focuses on audio-visual alignment.
The compressed latent feature $\mathcal{F}\in\mathbb{R}^{F\times H\times W\times C}$ is processed through non-overlapping patches, where spatial-temporal positional embeddings (Vaswani, 2017) preserve structural relationships. This compact representation enables efficient transformer processing while reducing memory requirements by 4$\times$ compared to conventional approaches.
Interleaved Fusion with Linear Attention: Given a sequence of latent features $\bm{Z}\in\mathbb{R}^{F\times P\times d}$, we employ a transformer architecture to predict the noise for removal. While the original DiT (Peebles and Xie, 2023) processes individual frames independently, our video generation framework extends this through interleaved spatiotemporal fusion. The naive approach of flattening $\bm{Z}$ into $(F\times P,d)$ tokens with standard attention incurs prohibitive quadratic complexity $\mathcal{O}(F^{2}P^{2})$ for high-resolution video generation.
Our architecture employs a dual-block decomposition strategy that alternates between spatial and temporal modeling, significantly reducing computational complexity from $\mathcal{O}(F^{2}P^{2})$ to $\mathcal{O}(F^{2}+P^{2})$. To reduce computational complexity, we replace the traditional softmax attention with ReLU linear attention (Katharopoulos et al., 2020). For spatial modeling, the input is first reshaped into $(F,P,d)$ dimensions to compute intra-frame attention using:
(1) $\bm{Z_{s}}=\text{LinearAttn}(\bm{Q}_{s},\bm{K}_{s},\bm{V}_{s})=\frac{\sum_{j=1}^{N}\phi(\bm{V}_{j})\phi(\bm{K}_{j})^{\top}\bm{Q}_{s}}{\sum_{j=1}^{N}\phi(\bm{K}_{j})^{\top}\bm{Q}_{s}}$
where $\bm{Q}_{s},\bm{K}_{s},\bm{V}_{s}\in\mathbb{R}^{F\times P\times d}$, $\phi(x)=\text{ReLU}(x)$. The linear attention mechanism further optimizes this by reusing shared terms across all queries: $\sum_{j=1}^{N}\text{ReLU}(\bm{K}_{j})\bm{V}_{j}^{T}\in\mathbb{R}^{d\times d}$.
For temporal modeling, the tensor to $(P,F,d)$ dimensions. Since $F\ll P$, we preserve the expressive power of standard softmax attention for inter-frame interactions: $\bm{Z_{t}}=\text{Attn}(\bm{Q}_{t},\bm{K}_{t},\bm{V}_{t})=\frac{\bm{V}_{t}(\bm{K}_{t}^{\top}\bm{Q}_{t})}{\sqrt{d}}$. Through this hybrid decomposition with linear attention, the overall complexity is reduced to $\mathcal{O}(F^{2}+Pd))$ achieving near-linear efficiency while preserving global receptive fields.
Mix-FFN block In spatial modeling, linear attention converges slower despite matching final performance. Following SANA (Xie et al., 2024a), we improve training efficiency by replacing the original MLP-FFN with Mix-FFN, which combines an inverted residual block, 3×3 depth-wise convolution, and GLU. The depth-wise convolution enhances local feature capture, compensating for ReLU linear attention’s weaker locality. Our experiments demonstrate this architecture achieves superior computational efficiency without compromising visual quality or temporal consistency compared to quadratic-complexity baselines in Fig. 1.
Figure 3. Multimodal fusion schemes: (a) Direct Fusion injects conditions via cross-attention modules; (b) Siamese Fusion uses parallel transformer for feature guidance; (c) Symbiotic Fusion achieves fusion through input concatenation and self-attention. The backbone architecture (left-side blocks) remains consistent across all approaches.
Schemes Integration State Params Modality Adapt.
Direct Cross-attention Static Dedicated Weak alignment
Siamese Self-attention/Add Dynamic Dedicated Flexible
Symbiotic Self-attention Dynamic Shared Strong
Table 1. Comparative analysis of three fusion schemes across three key dimensions: integration methods, condition states, and parameters sharing implementation
3.2. Analysis of Multimodal Fusion Schemes
After defining the transformer backbone, the next step is to propose a fusion scheme that exploits the input reference image and audio to achieve the multimodality-guided video generation task. Given the condition modality, we summarize the existing fusion solutions into the following three schemes based on the transformer paradigm:
  • Direct Fusion directly feeds the condition modality into each block of the backbone and fuses it with the inner features via cross-attention;
  • Siamese Fusion feeds the condition modality into another similar transformer, extracts features from each block, and fuses them with the features in the corresponding block of the backbone;
  • Symbiotic Fusion concatenates condition modality with inputs before feeding it into the backbone and achieving seamless fusion via inherent self-attention.
Integration methods: The choice between self-attention and cross-attention involves trade-offs in computational efficiency and optimization. Cross-attention, with smaller interaction matrices, is computationally lighter and easier to optimize, especially in limited data scenarios, as demonstrated in early text-to-image models (Rombach et al., 2022; Nichol et al., 2021). In contrast, self-attention, which requires larger matrices for multimodal token alignment, is more challenging to optimize but achieves higher embedding quality with sufficient data (Wang et al., 2024c). Alternatively, addition, which sums hidden states, provides a simpler and more computationally efficient approach, particularly useful when conditional inputs aim to highlight specific local features in the latent space.
Condition States: The static versus dynamic nature of conditions reflects their constraint strength. Dynamic states, as in Siamese and Symbiotic Fusion, enable hierarchical refinement of conditional features during training, aligning with findings from ReferenceNet (Hu, 2024) and ControlNet (Zhang et al., 2023b) that multi-level conditioning enhances generation controllability. This flexibility improves feature alignment and model performance.
Parameters Sharing: Siamese Fusion uses separate network parameters for conditions and latent noise, while Symbiotic Fusion shares the same parameters. Direct Attention, relying solely on cross-attention, also does not share parameters.
Modality Adaptation: The adaptability of fusion schemes to different modalities is key to their success in multimodal tasks. Direct Fusion, though offering weak alignment, is highly cost-effective for modalities with significant differences, such as text and image, making it ideal for resource-constrained scenarios. Siamese Fusion offers enhanced design flexibility by allowing adjustable modality alignment. This is exemplified in models such as ControlNet (Zhang et al., 2023b), which employs visual cue-level alignment, and AnimateAnyone (Hu, 2024), which utilizes pixel-level alignment. The strength of alignment can thus be precisely tailored to meet the specific requirements of the task at hand. Symbiotic Fusion enhances parameter efficiency compared to Siamese Fusion, achieving strong alignment with fewer parameters. This makes it particularly effective for large-scale applications, such as AnimateAnyone (Hu, 2024), where shared parameters ensure robust and efficient multimodal integration. The comparative analysis of the three fusion approaches are summarized in Tab. 1
3.3. Guidance on Reference Portrait and Audio
Given the reference image, we require that the character of the generated video is consistent with the reference. Since the image modality is close to the video, Symbiotic Fusion proves appropriate. Specifically, we first encode the reference image and obtain the portrait feature $\bm{P}\in\mathbb{R}^{P\times d}$ in the latent space. Then we temporally repeat the portrait feature $\bm{P}$ with $F$ frames and concatenate it with the latent feature $\bm{Z_{s}}\in\mathbb{R}^{F\times P\times d}$ along the spatial dimension, which can be formulated as: $\bm{Z}=\text{concat}\left(\,\bm{Z_{s}},\,\tilde{\bm{P}}\,\right)$ denotes the temporally extended portrait feature. This operation yields a concatenated latent feature $\bm{Z}\in\mathbb{R}^{F\times 2P\times d}$. By leveraging shared self-attention in the fused feature space, this method maintains consistent character representation across frames. In contrast, Direct Fusion uses shallow portrait features, leading to misalignment in deeper layers. While Siamese Fusion mitigates this by maintaining parallel transformer branches, it increases computational complexity without significant advantages over Symbiotic Fusion.
As for the audio, we require it to reasonably drive the animation, especially the lip motion (Zhang et al., 2023a). However, certain unconscious actions (e.g. blinking and head movements) are almost independent of the audio yet serve to diversity. Further considering that the modality of audio is quite different from the video, using Siamese Fusion and Symbiotic Fusion often causes ambiguity in the backbone learning, damaging the diversity. In a nutshell, Direct Fusion is concise yet effective for fusing the input audio. In addition, for the audio representation, to accommodate the tokenization mechanism of the transformer, we follow (Xu et al., 2024b) and extract richer semantic information from different layers of the wav2vec encoder (Schneider et al., 2019). Through nonlinear mapping, we obtain audio tokens $\bm{C}\in\mathbb{R}^{F\times T\times d^{\prime}}$ enriched with semantic content, which are aligned with latent features in the same feature space. Temporal-level audio representations are then synchronized with video sequences, resulting in the semantic audio representation. The experiments in Sec. 4.4 verify the efficacy of our mechanism in feature fusion.
Figure 4. The illustration of the long-duration generation.
3.4. Long-Duration Generation
Unlike the previous frame-by-frame generation approach (Shen et al., 2023; Stypułkowski et al., 2024), our end-to-end diffusion transformer generates a video clip in a single sampling process. In diffusion-based video generation tasks, ensuring proper temporal alignment is critical for maintaining coherence and consistency across generated video sequences. Additionally, due to the enhanced realism and dynamism of the talking videos we generate, direct transitions between consecutive clips are often not feasible. To address this, we introduce a memory bank that stores information from previously generated clips. This data undergoes patchification and temporal transpose operations, transforming it into $\bm{M}\in\mathbb{R}^{P\times f\times d}$, where $f$ represents the number of frames selected from the prior clips. Leveraging the scalability of the Transformer architecture, the memory bank directly provides prior information through Symbiotic Fusion with $\bm{Z_{t}}$. However, the inherent distribution discrepancy between training (real data) and inference (synthetic data) poses a fundamental challenge for diffusion models, with error accumulation becoming increasingly severe as video length grows.
To mitigate this latent risk, we develop a noise-regularized training strategy. Specifically, we first establish that diffusion models inherently retain residual noise even after finite-step denoising due to the discrete approximation of continuous noise schedules, which progressively accumulates as the video sequence extends. Building on the insight, we introduce a bounded stochastic noiser that perturbs the memory bank contents with controlled Gaussian noise $\epsilon\sim\mathcal{N}(0,{\sigma_{mini}}^{2}\mathbf{I})$ (see Fig. 4) into the memory bank during training. Inspired by Hallo2 (Cui et al., 2024), we further apply a patch-drop data augmentation strategy to the memory bank. Unlike Hallo2 which protects face patches from dropping, we employ completely random dropping to enhance the model’s robustness to the memory bank. This deliberate corruption serves two purposes: (1) it faithfully simulates inference-time distribution shifts, and (2) it preserves crucial temporal dependencies. The resulting framework demonstrates significantly improved robustness against error propagation in long-duration generation tasks.
Figure 5. The qualitative comparisons with other cutting-edge methods on the HDTF dataset. Our method achieves better audio-animation alignment (e.g. lip motions) and produces expressive results.
4. Experiment
4.1. Experimental Setups
Datasets We primarily conduct comprehensive experiments on two public datasets: HDTF (Zhang et al., 2021) and CelebV-HQ (Zhu et al., 2022). These datasets include videos that consist of individuals of various ages, ethnicities, and genders, captured in half-body or close-up shots against varied indoor and outdoor backgrounds. To ensure high-quality training data, we adopt the preprocess scheme used in (Xu et al., 2024b). We first exclude videos with scene changes, significant camera movements, excessive facial motion, and side-facing shots. Then we refine the remaining videos and retain only single-person speaking videos with strong audio-lip correspondence. After refinement, we obtain 356 clips with 13.95 hours from the HDTF, and 24438 clips with 44.91 hours from the CelebV-HQ dataset. During training, we extract 16-frame video clips from these datasets using a specific sampling interval, each frame is resized to $256\times 256$ resolution.
Evaluation Metrics The evaluation metrics utilized in the portrait image animation approach include Fréchet Inception Distance (FID), Fréchet Video Distance (FVD), Structural Similarity Index Measure (SSIM) and Expression-FID (E-FID). Specifically, FID and FVD measure the similarity between generated images and real data, with lower values indicating better performance and thus more realistic outputs (Karras et al., 2019). SSIM indices measure the structural similarity between ground truth videos and generated videos. E-FID leverages Inception features to assess image authenticity, providing a more precise fidelity measure. It extracts facial expression parameters via a reconstruction method (Deng et al., 2019) and computes FID to quantify the gap between generated and ground truth expressions.
Baseline In experiments, we compare our method with the publicly available implementations of DreamTalk (34M) (Ma et al., 2023), SadTalker (138M) (Zhang et al., 2023a), AniPortrait (2.8B) (Wei et al., 2024), Hallo (2.4B) (Xu et al., 2024b), Hallo (2.4B) (Cui et al., 2024) and EchoMimic (2.6B) (Chen et al., 2024). The evaluation is performed on the HDTF and CelebV-HQ Dataset, utilizing a training and testing split where $90\%$ of the identity data was allocated for training purposes.
Figure 6. The qualitative comparisons with the existing portrait image animation approaches on the CelebV-HQ dataset. Our method achieves better portrait consistency.
Method FID $\downarrow$ FVD $\downarrow$ SSIM $\uparrow$ E-FID $\downarrow$
DreamTalk 105.434 529.510 0.406 3.466
SadTalker 102.371 362.634 0.414 2.168
AniPortrait 78.284 331.117 0.623 2.903
EchoMimic 55.791 299.343 0.639 2.493
Hallo 45.754 285.988 0.698 2.235
Hallo2 44.343 241.628 0.657 2.017
Ours (Base) 46.240 289.385 0.666 2.587
Ours (Large) 31.761 219.119 0.632 1.962
Table 2. The quantitative comparisons with the existed portrait image animation approaches on the HDTF dataset.
Implementation details Our approach, in contrast to previous methods (Tian et al., 2024; Xu et al., 2024b; Wei et al., 2024) that rely on a two-stage training process, adopts an end-to-end training strategy. We set up models with two sizes: the Base model, which is suitable for simple fine-tuning, and the Large model, which demonstrates better generalization ability.
4.2. Quantitative Results
The comprehensive quantitative comparisons of HDTF and CelebV datasets are listed in Tab. 2 and Tab. 3, respectively. The results show that our method significantly outperforms other state-of-the-art methods in almost all metrics, achieving the best portrait consistency and audio-animation alignment. Notably, our approach achieves the lowest FID (31.761) and FVD (219.119) on the HDTF dataset, as well as the lowest FID (7.610) and FVD (33.175) on the CelebV-HQ dataset. These results show that our method generates both realistic and temporally consistent talking-head videos. For a fair comparison, each method generates the first video clip for every validation video, which is then resized to the same resolution for evaluation.
4.3. Qualitative Results
We have visualized the results of different methods on the HDTF dataset (Fig. 8) and the CelebV-HQ dataset (Fig. 9). From the visualization results, it is evident that our method performs exceptionally well on both datasets. In comparison to the visual cues most strongly associated with audio (such as facial muscle and lip motions), our results closely align with the standard Ground Truth. Benefitting from our Symbiotic Fusion, our generated videos maintain strict portrait consistency. Furthermore, by employing Direct Fusion of audio, the characters exhibit strong expressiveness and diversity in our generated videos while achieving audio-animation alignment.
Method FID $\downarrow$ FVD $\downarrow$ SSIM $\uparrow$ E-FID $\downarrow$
DreamTalk 64.297 293.905 0.448 5.915
SadTalker 42.017 77.802 0.469 3.372
AniPortrait 33.408 107.236 0.578 3.722
EchoMimic 12.932 59.944 0.567 3.671
Hallo 8.504 46.431 0.580 3.280
Hallo2 8.734 41.213 0.543 3.128
Ours (Base) 9.585 36.159 0.598 3.349
Ours (Large) 7.610 33.175 0.584 3.038
Table 3. The quantitative comparisons with the existed portrait image animation approaches on the CelebV-HQ dataset.
Reference Portrait FID $\downarrow$ FVD $\downarrow$ SSIM $\uparrow$ E-FID $\downarrow$
Direct Fusion 52.999 389.01 0.497 3.149
Siamese Fusion 47.779 286.692 0.651 2.597
Symbiotic Fusion 46.240 289.385 0.666 2.587
Table 4. Experimental results of three multimodal fusion schemes, guided by reference portraits, on the HDTF validation dataset, employing the Letstalk-B model.
Driven Audio FID $\downarrow$ FVD $\downarrow$ SSIM $\uparrow$ E-FID $\downarrow$
Symbiotic Fusion 51.880 403.628 0.471 3.398
Siamese Fusion 48.041 318.637 0.585 2.706
Direct Fusion 46.240 289.385 0.666 2.587
Table 5. Experimental results of three multimodal fusion schemes, driven by audio inputs, on the HDTF validation dataset, employing the Letstalk-B model.
4.4. Ablation Study
In this section, we analyze the impact of different fusion schemes as illustrated in Fig. 3 on guidance using reference portrait and audio. Specifically, for analyzing Siamese Fusion, we have to design Siamese transformers for both reference portrait and audio to produce guidance for each block of the backbone. For the reference portrait, we leverage a DiT model pre-trained on ImageNet. For the audio, we define a transformer whose blocks are equipped with two-layer MLPs. The parameters of these transformers are simultaneously updated during training.
Analysis on the Reference Portrait Fusion Portrait consistency is crucial to the portrait image animation task and highly relevant to the portrait fusion scheme used. We evaluate the synthesis quality of portrait fusion using three schemes, and the quantitative results are shown in Tab. 4. Symbiotic Fusion almost surpasses the other two schemes, realizing better portrait consistency and video quality. Besides, while Siamese Fusion produces comparable results, it introduces additional parameters (i.e. siamese transformer) and complicates the learning task.
Analysis on the Driven Audio Fusion The way we fuse audio affects both audio-animation alignment and diversity. Audio-animation alignment refers to specific visual cues (Zhang et al., 2023a) (e.g. facial muscle and lip motions), whereas the diversity is related to the speech-agnostic motions (e.g. blinking and neck moving) that often appear as unconscious actions. We observe that diversity mainly comes from the transformer backbone during denoising. Thus, we aim to fuse the audio to achieve alignment while preserving the backbone’s inherent diversity. Quantitative results in Tab. 5 indicate that Direct Fusion outperforms the other schemes in establishing the correspondence between audio and animation.
Method FID $\downarrow$ FVD $\downarrow$
Baseline model with Memory Bank 52.243 276.16
+ Regularized Training (w/o patch drop) 32.746 264.709
+ Regularized Training (w patch drop) 34.821 251.684
Table 6. Experimental results of long video generation. Experimental results comparing three approaches for 64-second video generation (100 clips): (1) memory bank baseline, (2) regularized training, and (3) patch drop augmentation.
4.5. Long-Duration Generation
To validate the effectiveness of our proposed approach for long video generation, we conducted experiments to analyze the quality of 100 consecutive video clips (about 64 second) generated under three settings: (1) the baseline model (LetsTalk-L) with with the memory bank, (2) the method augmented with regularized training strategy, and (3) the method further enhanced with the patch drop augmentation. Our results in Tab. 6 demonstrate that the memory bank significantly improves temporal consistency, while the regularized training strategy further mitigates degradation in later clips, leading to higher overall video quality. The results demonstrate that regularized training strategy primarily enhances single-frame quality, significantly reducing FID from 52.243 to 32.746, while maintaining temporal consistency (FVD 264.709). On the other hand, patch drop augmentation further improves long-term motion coherence, achieving the best FVD (251.684) with only a slight FID increase (34.821), confirming its complementary role in stabilizing long-duration generation.
5. Conclusion
We introduced LetsTalk, an efficient linear diffusion transformer for high-fidelity talking head synthesis, tackling core challenges in multimodal fusion and long-term temporal coherence. Our approach integrates a deep compression autoencoder for compact latent representation, alongside a systematic exploration of fusion strategies—yielding an optimal framework with Symbiotic Fusion for portrait consistency and Direct Fusion for audio-video synchronization. The memory bank mechanism and noise-regularized training further mitigate error accumulation in extended sequences. Extensive experiments demonstrate LetsTalk’s state-of-the-art performance, generating lifelike, temporally coherent animations with enhanced diversity, all while maintaining superior efficiency through its optimized architecture.
Appendix A Preliminaries
Diffusion Formulation The diffusion process incrementally corrupts data $x_{0}$ with Gaussian noise: $q(x_{t}|x_{0})=\mathcal{N}(x_{t};\sqrt{\bar{\alpha}_{t}}x_{0},(1-\bar{\alpha}_{t})\mathbf{I})$, where $\bar{\alpha}_{t}$ are predetermined hyperparameters. By reparameterizing, we obtain: $x_{t}=\sqrt{\bar{\alpha}_{t}}x_{0}+\sqrt{1-\bar{\alpha}_{t}}\epsilon_{t}$, where $\epsilon\sim\mathcal{N}(0,\mathbf{I})$, forming a Markov chain over $T$ steps.
Diffusion models are trained to learn the reverse process, modeled as: $p_{\theta}(x_{t-1}|x_{t})=\mathcal{N}(\mu_{\theta}(x_{t}),\Sigma_{\theta}(x_{t}))$, with neural networks parameterizing $p_{\theta}$. This is optimized via the variational lower bound of $\log p(x_{0})$:
(2) $\begin{split}\mathcal{L}(\theta)=&\sum_{t}\mathcal{D}_{\text{KL}}(q^{*}(x_{t-1}|x_{t},x_{0})\,\|\,p_{\theta}(x_{t-1}|x_{t}))\\ &-p(x_{0}|x_{1}).\end{split}$
With both $q$ and $p_{\theta}$ Gaussian, $D_{\text{KL}}$ is computed from their means and covariances. Reparameterizing $\mu_{\theta}$ as a noise-prediction network $\epsilon_{\theta}$ enables training via mean-squared error between $\epsilon_{\theta}(x)$ and sampled noise $\epsilon_{t}$:
(3) $\mathcal{L}_{\text{simple}}(\theta)=\|\epsilon_{\theta}(x)-\epsilon(x)\|_{2}^{2}.$
To incorporate a learnable reverse covariance $\Sigma_{\theta}$, we minimize the full $D_{\text{KL}}$ term, following Nichol and Dhariwal: optimizing $\theta$ with $\mathcal{L}_{\text{simple}}$ and $\Sigma_{\theta}$ with the full $\mathcal{L}$.
Diffusion Transformer To better handle the multi-modality fusion in the talking head task, we are motivated to extend the Diffusion Transformer (DiT) in our framework. DiT builds upon the Vision Transformer (ViT) architecture, which processes sequences of image patches. This sequential patch-based processing mechanism natively supports the fusion of multi-modality sufficiently and efficiently. Different from ViT, DiT incorporates a variational autoencoder (VAE) (Kingma, 2013) in the framework, thus runs in the latent space, rather than pixel space. Such design enables DiT to handle the expensive high-resolution generation task, which brings more obvious benefits in the video generation task.
Model Patch size Layer num N Hidden size D Paramaters
LetsTalk-B 1 12 768  [^2] M
LetsTalk-L 1 24 1024 773
Table 7. Details of LetsTalk models. We follow ViT and DiT model configurations for different model sizes.
Appendix B Network Architecture Details
B.1. Backbone
In LetsTalk, we borrow the Variational AutoEncoder (VAE) (Kingma, 2013) from Stable Diffusion (Rombach et al., 2022). Given the reference portrait image and noisy video, we first encode them using the encoder in VAE. Then we patchify and flatten them along the spatial dimension. Finally, we use an MLP to project them into the latent space. The embeddings in the latent space will be fed into our transformer backbone. Specifically, our LetsTalk model is constructed using a series of N transformer blocks, each with a hidden dimension of D and H multi-head attention heads. Inspired by ViT and DiT, we explore two configurations of LetsTalk with varying numbers of parameters, while maintaining a fixed patch size of 2, as detailed in  Tab. 7. Via the encoder in VAE, the downsampling factor is $8$, the input reference portrait image $\bm{I}_{\text{p}}\in\mathbb{R}^{256\times 256\times 3}$ and the noisy video $\bm{V}\in\mathbb{R}^{16\times 256\times 256\times 3}$ are encoded as a portrait feature $\bm{P}\in\mathbb{R}^{32\times 32\times 4}$ and a video feature $\bm{F}\in\mathbb{R}^{16\times 32\times 32\times 4}$. Next, the two features are partitioned according to $2\times 2$ patch size and flattened as $\bm{P}\in\mathbb{R}^{256\times 16}$ and $\bm{F}\in\mathbb{R}^{16\times 256\times 16}$. Take our base model (i.e. LetsTalk-Base) as an example, we use an MLP to project both features into the latent space as $\bm{P}\in\mathbb{R}^{256\times 768}$ and $\bm{F}\in\mathbb{R}^{16\times 256\times 768}$.
B.2. Adaptive Integration
Following VDT(Lu et al., 2023), we insert an adaptive full-connected layer between attention layers:
(4) $\bm{Z}_{s}=\bm{Z_{s}}+\alpha\cdot\text{FC}(\text{Temp-Attn}(\bm{Z_{t}}))$
where $\alpha$ is a learnable scalar and $\text{FC}(\cdot)$ denotes a fully-connected layer. This enables adaptive fusion of temporal dynamics into spatial representations.
B.3. Audio Encoder
We resample the audio signal to a 16 KHz sample rate. For audio feature extraction, we use wav2vec (Schneider et al., 2019), employing embeddings from the last 12 layers to capture a richer range of semantic information across different levels of the audio representation. To ensure frame-by-frame alignment between audio and animation, we match the audio features with corresponding frames of the video clips. Additionally, we use a sliding window (multiple audio frames) for each video frame to maintain audio continuity and preserve more accurate audio information. To extract audio representations suitable for transformer tokenization, we designed an audio projection module. This module first flattens the extracted audio features and then selects a fixed number of context tokens (32 in our experiments). Simultaneously, the audio representations are mapped to the same dimensional feature space as the latent noise.
B.4. Siamese Transformer
The Siamese Fusion scheme has been adopted in AnimateAnyone (Hu, 2024), EMO (Tian et al., 2024), Hallo (Xu et al., 2024b), and so on. The Siamese Fusion scheme aims to update a backbone-like network simultaneously and extract the features to fuse it into the backbone block-wise. Since these methods follow Stable Diffusion and use a U-Net to predict the noise for denoising, they adopt a similar U-Net as the siamese network. Considering LetsTalk as a Diffusion Transformer, we introduce a ViT (Dosovitskiy, 2020) pre-trained on ImageNet to achieve the Siamese Fusion scheme for the reference portrait image. For the input audio, we propose a transformer whose blocks consist of a two-layer MLP to implement the Siamese Fusion scheme.
Figure 7. Portrait animation results without audio guidance for three fusion schemes. Siamese and Symbiotic fusion schemes with high fusion compactness limit the diversity of video generation, while Direct fusion preserves this diversity.
B.5. Impact of Audio-guidance on Diversity
To validate the impact of different fusion schemes on diversity, we serve LetsTalk as an unconditional video generator [^1] Filling audio embeddings with 00 for fusion. and compare the diversity of character actions in the generated videos. The results in Fig. 7 demonstrate that the deeper the fusion compactness, the more static the generated video. The tightest Symbiotic Fusion generates a portrait with almost no motions, Siamese Fusion produces slight motions, while the least tight Direct Fusion we adopt results in natural motions.
Appendix C Additional Implementation Details
C.1. Diffusion
We leverage a pre-trained variational autoencoder model from Stable Diffusion 1.4, with all our diffusion models operating in the latent space of this VAE. After sampling a latent from the diffusion model, we decode it to pixel space using the VAE decoder. We adopt the hyperparameters from ADM (Dhariwal and Nichol, 2021), specifically employing a linear variance schedule with $t_{\text{max}}=1000$, ranging from $1\times 10^{-4}$ to $2\times 10^{-2}$. Additionally, we use ADM’s parameterization of the covariance $\Sigma_{\theta}$ and their approach for embedding input timesteps.
C.2. Training
We initialize LetsTalk from a pre-trained Latte (Ma et al., 2024) model on the FaceForensics dataset (Rössler et al., 2018). Following the training strategy outlined in Latte, we adopt gradient clipping to stabilize training and prevent gradient explosion. For training, we use the AdamW optimizer with a constant learning rate of $1\times 10^{-4}$ and no weight decay. The batch size ranges from 10 to 16, and the only data augmentation applied is horizontal flips. Each model is trained for 300k to 500k iterations. Following common practices in generative modeling works (Peebles and Xie, 2023; Bao et al., 2023), an exponential moving average (EMA) weight with a decay rate of $0.9999$ is maintained throughout training, and all reported results are directly derived from the EMA model. Additionally, we utilize a pre-trained VAE from Stable Diffusion 1.4, which remains frozen during training. Training is conducted on both RTX 3090 and A800 GPUs, using automatic mixed precision to optimize both performance and memory usage.
C.3. Inference
During the inference stage, the network takes a single reference portrait and driving audio as input, generating a video sequence that animates the reference image based on the corresponding audio. The length and resolution of the generated video clip during inference need to remain consistent with those used during training, which in our experiments are $16$ frames and a resolution of $256\times 256$. We use DDPM for the sampling strategy, with a total of $250$ sampling steps.
C.4. Diffusion
We leverage a pre-trained variational autoencoder model from Stable Diffusion 1.4, with all our diffusion models operating in the latent space of this VAE. After sampling a latent from the diffusion model, we decode it to pixel space using the VAE decoder. We adopt the hyperparameters from ADM (Dhariwal and Nichol, 2021), specifically employing a linear variance schedule with $t_{\text{max}}=1000$, ranging from $1\times 10^{-4}$ to $2\times 10^{-2}$. Additionally, we use ADM’s parameterization of the covariance $\Sigma_{\theta}$ and their approach for embedding input timesteps.
Appendix D Qualitative Comparisons
In this section, we report comprehensive qualitative comparisons of LetsTalk with other cutting-edge methods and demonstrate more applications of LetsTalk. We show more results on the HDTF (Zhang et al., 2021) and Celeb-HQ (Zhu et al., 2022) benchmarks ( Fig. 8 and Fig. 9).
We show the portrait animation results of LetsTalk driven by English (Fig. 10), Chinese (Fig. 11), and singing audios (Fig. 12). We further use the powerful AI assistant (Doubao) to generate portraits for comparisons (Fig. 13). Note that LetsTalk is only trained on the HDTF and Celeb-HQ datasets, which only contain daily conversations of real persons.
Figure 8. The qualitative comparisons with other cutting-edge methods on the HDTF dataset.
Figure 9. The qualitative comparisons with other cutting-edge methods on the Celeb-HQ dataset.
Appendix E Visualization
We provide video visualizations in our attachment. Notably, the model used to generate the video is only trained on HDTF (Zhang et al., 2021) and CelebV-HQ (Zhu et al., 2022), which only include English conversation scenarios of real portraits. In addition, we use Chinese audio and singing audio to drive real portraits. Furthermore, LetsTalk can drive AI-generated portraits [^2] We use Doubao to generate AI-generated portraits in the attached video.. The visualization results demonstrate that LetsTalk has great generalization and achieves remarkable results outside the dataset.
Figure 10. English audio-driven results of LetsTalk. Note that each column corresponds to the same audio.
Figure 11. Chinese audio-driven results of LetsTalk. Note that each column corresponds to the same audio.
Figure 12. Singing audio-driven results of LetsTalk. Note that each column corresponds to the same audio.
Figure 13. Results of LetsTalk driving AI-generated portraits using English audio. Note that each column corresponds to the same audio.