### v3

# Efficient Long-duration Talking Video Synthesis with Linear Diffusion Transformer under Multimodal Guidance

Abstract
[…]
Index Terms:
Talking Head Generation, Long-duration Generation, Multimodal Guidance, Linear Diffusion Transformer, Deep Compression Autoencoder
[…]
I Introduction
Animating a portrait image with speech, also known as talking head generation, is a challenging task that aims to generate a speaking portrait from a single image with audio. Naturally driving the reference portrait according to the given audio shows great potential in video gaming, film, digital marketing, etc. Despite recent progress, long-duration talking video synthesis still faces persistent challenges in simultaneously achieving high video quality, portrait and temporal consistency, and computational efficiency. As video length increases, issues such as visual degradation, loss of identity consistency, temporal incoherence, and error accumulation become increasingly prominent, severely impacting the realism and reliability of generated results.
Previous works [1, 2] use a 3D morphable model (3DMM) [3, 4, 5] to represent reference portraits and drive them with audio signals. These approaches input driven 3DMM parameters (e.g., coefficients, poses) into neural renderers, achieving temporal and portrait consistency yet lacking output diversity, often showing static features like unblinking eyes and rigid hair. In contrast, audio-to-2D facial motion methods [6, 7, 8, 9, 10] directly align audio with facial dynamics via latent space learning. However, their limited disentanglement of static and dynamic components compromises both temporal consistency and animation diversity.
Recently, diffusion models [11, 12, 13] have emerged as a powerful paradigm for generating realistic 2D images [13, 14] and have demonstrated remarkable success in video generation [15, 16]. Notably, AnimateAnyone [17] introduces motion modules and ReferenceNet into diffusion models to produce realistic and pose-controllable human animations, showcasing the capability of end-to-end diffusion models to effectively capture the correspondence between conditional signals and human motion. Building on this, recent methods [18, 19, 20] have extended this approach by replacing pose signals with audio signals for talking head generation. These methods leverage diffusion models to generate diverse unconscious actions, significantly enhancing the richness of generated content.
However, several critical challenges remain unaddressed, particularly in the context of long-duration video generation. As the length of generated videos increases, existing methods struggle to maintain high visual quality, portrait consistency, and temporal coherence. Progressive accumulation of residual noise and artifacts (e.g., facial distortions or flickering) becomes increasingly severe over extended sequences, ultimately degrading output fidelity and realism. In addition to these long-duration challenges, current approaches also suffer from suboptimal efficiency in audio-driven animation, primarily due to: (1) reliance on auxiliary spatial signals (e.g., predefined movement regions) as a crutch for limited audio controllability, which sacrifices motion diversity and naturalness; (2) redundant parameter overhead from fine-tuning generic text-to-image/video models, despite their misalignment with animation-specific tasks; and (3) inefficient architectures, exemplified by low-compression-rate autoencoders and vanilla attention mechanisms with high computational complexity. Furthermore, despite the focus on conditional video generation, there is a lack of systematic exploration into the intrinsic modeling principles for effectively integrating multimodal conditions, such as reference portraits and driving audio.
In this work, we introduce LetsTalk, a diffusion transformer framework designed for efficient and high-quality generation of realistic, long-duration talking videos under multimodal conditions. A key component for long video generation in LetsTalk is the memory bank mechanism, which enables each generated clip to access information from previous clips, ensuring smooth transitions and contextual continuity. To address the discrepancy between training and inference, we propose a noise-regularized training strategy that explicitly simulates inference-time noise accumulation, thereby mitigating error accumulation and sampling artifacts as the video extends.
[…]
Regarding multimodal guidance, we define it in this task as the fusion of video generation with reference portrait and driving audio. To better exploit multimodality for guiding long-duration video generation, we first summarize three fusion schemes from shallow to deep according to the compactness of the fusion: 1) Direct Fusion, 2) Siamese Fusion, and 3) Symbiotic Fusion. We analyze these fusion schemes in terms of integration, conditional state, and parameter sharing, systematically summarizing their adaptability to different modalities. We further investigate the underlying factors influencing such adaptability, thereby bridging theoretical gaps in conditional guidance modeling. To effectively guide long-duration video generation, we use Symbiotic Fusion for the reference image, ensuring strong portrait consistency throughout extended sequences due to its visual similarity to the video. For audio, we adopt Direct Fusion to align animations (e.g., lip motion, expressions) with the audio while maintaining diversity in movements (e.g., blinking, head motion) over long video durations.
We summarize the contributions of this work as follows
  • We propose LetsTalk, a novel diffusion transformer framework for long-duration talking video generation, which simultaneously achieves high video quality, portrait and temporal consistency, and computational efficiency.
  • We introduce a memory bank mechanism combined with a noise-regularized training strategy, which explicitly maintains contextual continuity and effectively mitigates error accumulation and sampling artifacts during long video generation.
  • We design an efficient spatiotemporal modeling architecture, leveraging a deep compression autoencoder and linear attention, which achieves comparable visual fidelity with only 1/8 the parameters of previous methods.
  • We systematically analyze multimodal fusion schemes and adopt deep Symbiotic Fusion for portrait features and shallow Direct Fusion for audio, enabling LetsTalk to achieve state-of-the-art visual quality and temporal coherence in talking video generation.
II Related Work
II-A Diffusion Models for Video Generation
Diffusion models have demonstrated exceptional potential across various domains, particularly in image [11, 12] and video generation [22, 17, 23, 24]. A notable example is Stable Diffusion [13], which utilizes a UNet architecture to iteratively produce high-quality images from text prompts after extensive training on large-scale text-image datasets [25]. This approach has been extended to more complex tasks like video generation [22, 17], where models such as Video Diffusion Models (VDM)[23] apply the diffusion process with space-time factorized training, and ImagenVideo[26] enhances VDM with cascaded diffusion models for high-resolution outputs. Further advances, including VideoComposer [27], AnimateDiff [22], and VideoCrafter [28], focus on synthesizing videos by conditioning the diffusion process on images and blending image latents. A significant direction in the evolution of diffusion models is the incorporation of patch sequences of images and transformers, exemplified by the Diffusion Transformer (DiT) architecture [29], which facilitates scaling to larger datasets and model parameters. Recent works, such as Latte [15] and VDT [16], have extended the DiT architecture for video generation by incorporating temporal mechanisms, achieving strong results by training diffusion models from scratch. Building on this, our study also adopts an end-to-end approach for high-fidelity, temporally consistent video generation.
II-B Audio-driven Portrait Image Animation
In recent years, the field of portrait image animation has advanced significantly, especially in generating realistic and expressive talking head animations from static images paired with audio inputs. Early efforts primarily focused on improving lip synchronization, with models like LipSyncExpert [7] achieving high accuracy for specific identities. Subsequently, methods such as ATVGnet [6] and MakeItTalk [8] introduced facial landmarks and disentangled representations to improve video frame generation. The integration of 3D information in models like SadTalker [1], Genface [30], and DiffTalk [31] further improved head movements and expressions. A significant advancement came with the introduction of diffusion models, including Diffused Heads [32], DreamTalk [33], and VividTalk [34], which pushed the field forward by producing high-fidelity, natural talking head videos across diverse identities, focusing on improved lip synchronization and varied speaking styles. Recent innovations like VASA-1 [35], EMO [36] and AniPortrait [2] have focused on creating emotionally expressive talking faces, capturing a wide range of facial nuances and movements. AniTalker [37] introduced a universal motion model for realistic facial movements with minimal labeled data, enabling dynamic avatars. LivePortrait [38] extends the implicit-keypoint framework, balancing efficiency and controllability. Most recently, Hallo [18, 20], EchoMimic [19] and X-Portrait [39] have adopted the latent diffusion framework pioneered in AnimateAnyone [17], adding more control over elements such as lip movements, expressions, and poses.
II-C Conditionally Guided Diffusion Model
The impressive generative capabilities of diffusion models have made the integration of control information during generation a crucial area of research. Text-guided control methods focus on refining prompts, manipulating CLIP features, and modifying cross-attention mechanisms [40, 41, 42, 43, 44]. ControlNet [14] enhances large pretrained text-to-image diffusion models by incorporating spatially localized, task-specific conditioning. In video generation, AnimateAnyone [17] introduces the ReferenceNet to ensure the consistency of fine-grained appearance details from reference images. Gen-1 [45] integrates structural guidance through depth maps and cross-attention, while MCDiff [46], LaMD [47], and MotionCtrl [48] focus on motion-guided video generation, enhancing the realism of human actions. For talking video synthesis, previous approaches primarily relied on image diffusion models to generate dynamic portraits frame by frame, starting from reference portraits [31, 32]. Recent methods often incorporate ReferenceNet for processing reference portraits and cross-attention mechanisms for handling audio information [36, 18]. While control over both images and videos in latent diffusion models has seen significant advancements, the theoretical understanding of various conditioning strategies remains limited, particularly in handling multiple modalities (e.g., image, video, audio). To address this gap, we investigate conditioning strategies in diffusion transformers, aiming to improve control and generation quality across modalities.
III Methodology
Figure 2: The illustration of our method. LetsTalk is designed for robust long-duration talking video generation. It integrates a deep compression autoencoder to reduce spatial redundancy while preserving temporal features, and transformer blocks with temporal and spatial attention to capture intra-frame details and long-range temporal correspondences. After obtaining portrait and audio embeddings, Symbiotic Fusion is used for the portrait embedding and Direct Fusion for the audio embedding, providing effective multimodal guidance for long video synthesis. The portrait embedding is repeated along the frame axis to match the noise embedding for consistent conditioning. To support long video generation, we introduce a memory bank module to preserve temporal consistency across extended sequences. We also propose a noise-regularized training strategy to align the memory bank between training and inference, enabling stable and high-fidelity long-duration generation.
We present LetsTalk, a diffusion transformer for audio-driven portrait animation, as illustrated in Fig. 2. This section first introduces the diffusion formulation and transformer backbone in Section III-A. We then detail our long-range consistency mechanism, including the memory bank and noise-regularized training, in Section III-C. Next, we describe the efficient spatiotemporal transformer backbone in Section III-B. Finally, we analyze multimodal fusion schemes and our guidance strategy in Section III-D and Section III-E.
III-A Preliminaries
Diffusion Formulation The diffusion process incrementally corrupts data $x_{0}$ with Gaussian noise: $q(x_{t}|x_{0})=\mathcal{N}(x_{t};\sqrt{\bar{\alpha}_{t}}x_{0},(1-\bar{\alpha}_{t})\mathbf{I})$, where $\bar{\alpha}_{t}$ are predetermined hyperparameters. By reparameterizing, we obtain: $x_{t}=\sqrt{\bar{\alpha}_{t}}x_{0}+\sqrt{1-\bar{\alpha}_{t}}\epsilon_{t}$, where $\epsilon\sim\mathcal{N}(0,\mathbf{I})$, forming a Markov chain over $T$ steps.
Diffusion models are trained to learn the reverse process, modeled as: $p_{\theta}(x_{t-1}|x_{t})=\mathcal{N}(\mu_{\theta}(x_{t}),\Sigma_{\theta}(x_{t}))$, with neural networks parameterizing $p_{\theta}$. This is optimized via the variational lower bound of $\log p(x_{0})$:
$\begin{split}\mathcal{L}(\theta)=&\sum_{t}\mathcal{D}_{\text{KL}}(q^{*}(x_{t-1}|x_{t},x_{0})\,\|\,p_{\theta}(x_{t-1}|x_{t}))\\ &-p(x_{0}|x_{1}).\end{split}$ (1)
With both $q$ and $p_{\theta}$ Gaussian, $D_{\text{KL}}$ is computed from their means and covariances. Reparameterizing $\mu_{\theta}$ as a noise-prediction network $\epsilon_{\theta}$ enables training via mean-squared error between $\epsilon_{\theta}(x)$ and sampled noise $\epsilon_{t}$:
$\mathcal{L}_{\text{simple}}(\theta)=\|\epsilon_{\theta}(x)-\epsilon(x)\|_{2}^{2}.$ (2)
To incorporate a learnable reverse covariance $\Sigma_{\theta}$, we minimize the full $D_{\text{KL}}$ term, following Nichol and Dhariwal: optimizing $\theta$ with $\mathcal{L}_{\text{simple}}$ and $\Sigma_{\theta}$ with the full $\mathcal{L}$.
Diffusion Transformer To better handle the multi-modality fusion in the talking head task, we are motivated to extend the Diffusion Transformer (DiT) in our framework. DiT builds upon the Vision Transformer (ViT) architecture, which processes sequences of image patches. This sequential patch-based processing mechanism natively supports the fusion of multi-modality sufficiently and efficiently. Different from ViT, DiT incorporates a variational autoencoder (VAE) [49] in the framework, thus runs in the latent space, rather than pixel space. Such design enables DiT to handle the expensive high-resolution generation task, which brings more obvious benefits in the video generation task.
III-B Efficient SpatioTemporal Transformer with Linear Attention
Deep Compression Autoencoder: We build upon latent diffusion models and employ a deep compression autoencoder [21] that downsamples images by a factor of 32, maintaining reconstruction quality comparable to SDXL’s autoencoder. This design leverages the redundancy in high-resolution video frames and decouples compression from denoising, allowing the autoencoder to preserve crucial temporal features (e.g., lip-sync dynamics) while the diffusion model focuses on audio-visual alignment. Compared to conventional autoencoders, this approach significantly reduces memory and computation requirements while maintaining high fidelity.
[…]
Interleaved Fusion with Linear Attention: Given a sequence of latent features $\bm{Z}\in\mathbb{R}^{F\times P\times d}$, we employ a transformer architecture to predict the noise for removal. Unlike the original DiT [29], which processes each frame independently, our framework introduces interleaved spatiotemporal fusion to jointly model spatial and temporal dependencies.
A naive approach would flatten $\bm{Z}$ into $(F\times P,d)$ tokens and apply standard self-attention, resulting in prohibitive quadratic complexity $\mathcal{O}(F^{2}P^{2})$ for high-resolution videos. To address this, we adopt a dual-block decomposition strategy that alternates between spatial and temporal attention blocks, significantly reducing the computational cost to $\mathcal{O}(F^{2}+P^{2})$.
For spatial modeling, we reshape the input to $(F,P,d)$ and apply ReLU-based linear attention [51] within each frame:
$\bm{Z}_{s}=\text{LinearAttn}(\bm{Q}_{s},\bm{K}_{s},\bm{V}_{s})=\frac{\sum_{j=1}^{N}\phi(\bm{V}_{j})\phi(\bm{K}_{j})^{\top}\bm{Q}_{s}}{\sum_{j=1}^{N}\phi(\bm{K}_{j})^{\top}\bm{Q}_{s}},$ (3)
where $\phi(x)=\text{ReLU}(x)$ and $\bm{Q}_{s},\bm{K}_{s},\bm{V}_{s}\in\mathbb{R}^{F\times P\times d}$.
For temporal modeling, we transpose the tensor to $(P,F,d)$ and apply standard softmax attention across frames, allowing us to preserve the expressive power of full attention for inter-frame interactions.
By interleaving these spatial and temporal attention blocks, our architecture achieves near-linear efficiency while maintaining a global receptive field, enabling scalable and effective modeling of high-resolution, long-duration talking videos.
Mix-FFN block In spatial modeling, linear attention converges slower despite matching final performance. Following SANA [21], we improve training efficiency by replacing the original MLP-FFN with Mix-FFN, which combines an inverted residual block, 3×3 depth-wise convolution, and GLU. The depth-wise convolution enhances local feature capture, compensating for ReLU linear attention’s weaker locality. This enhances local feature capture and compensates for the weaker locality of linear attention.
Adaptive Integration Following VDT [16], we insert an adaptive fully-connected layer between attention layers to fuse temporal dynamics into spatial representations:
$\bm{Z}_{s}=\bm{Z_{s}}+\alpha\cdot\text{FC}(\text{Temp-Attn}(\bm{Z_{t}}))$ (4)
where $\alpha$ is a learnable scalar and $\text{FC}(\cdot)$ denotes a fully-connected layer. This enables adaptive fusion of temporal dynamics into spatial representations.
Figure 3: The illustration of the long-duration generation.
III-C Memory Bank and Noise-Regularized Training for Long-Term Consistency
In diffusion-based video generation, maintaining temporal alignment and consistency across long sequences is essential. However, as the video extends, the accumulation of residual noise and sampling artifacts can threaten temporal coherence and visual quality. This is further exacerbated by the mismatch between training and inference: during training, the memory bank is constructed from ground-truth frames, while during inference, it is updated with generated frames.
Memory Bank: To address these challenges, we introduce a memory bank that stores information from previously generated clips. Unlike previous frame-by-frame generation approaches [31, 32], our end-to-end diffusion transformer generates an entire video clip in a single sampling process. The memory bank is patchified and temporally transposed, resulting in $\bm{M}\in\mathbb{R}^{P\times f\times d}$, where $f$ denotes the number of frames selected from prior clips. The memory bank provides prior information to the current generation through Symbiotic Fusion with $\bm{Z_{t}}$, leveraging the scalability of the Transformer architecture. This design ensures smooth transitions and temporal consistency across video segments, and is especially advantageous for long-duration video generation.
Theoretical Motivation: We observed that DDPM-based sampling cannot fully remove noise due to two main factors:
  • Model prediction error: The predicted noise $\epsilon_{\theta}(x_{t},t)$ only approximates the true noise $\epsilon$, leaving some residual noise after denoising.
  • Sampling stochasticity: At each reverse step, new Gaussian noise is added:
    $x_{t-1}=\frac{1}{\sqrt{\alpha_{t}}}\left(x_{t}-\frac{1-\alpha_{t}}{\sqrt{1-\bar{\alpha}_{t}}}\epsilon_{\theta}(x_{t},t)\right)+\sigma_{t}z,\quad z\sim\mathcal{N}(0,I)$ (5)
    The term $\sigma_{t}z$ introduces inherent randomness.
Consequently, even after finite-step denoising, diffusion models retain residual noise, which accumulates as the video sequence extends. This motivates our noise-regularized memory bank, which differs from previous long video generation methods.
Noise-Regularized Training Strategy. To bridge the gap between training and inference, we propose a noise-regularized training strategy. We perturb the memory bank with controlled Gaussian noise, $\epsilon\sim\mathcal{N}(0,{\sigma_{mini}}^{2}\mathbf{I})$, simulating inference-time noise accumulation. Specifically, we randomly sample a diffusion timestep $t$ from the first 50 steps and add the corresponding noise to the memory bank during training. Additionally, inspired by Hallo2 [20], we apply patch-drop augmentation by randomly masking 25% of spatial patches in the memory bank, further enhancing model robustness.
This deliberate corruption enables the model to learn to correct for artifacts and maintain temporal dependencies. Our framework thus achieves robust and stable long-duration video synthesis, as demonstrated in Fig. 3.
Figure 4: Multimodal fusion schemes: (a) Direct Fusion injects conditions via cross-attention modules; (b) Siamese Fusion uses parallel transformer for feature guidance; (c) Symbiotic Fusion achieves fusion through input concatenation and self-attention. The backbone architecture (left-side blocks) remains consistent across all approaches.
Schemes Integration State Params Modality Adapt.
Direct Cross-attention Static Dedicated Weak alignment
Siamese Self-attention/Add Dynamic Dedicated Flexible
Symbiotic Self-attention Dynamic Shared Strong
TABLE I: Comparative analysis of three fusion schemes across three key dimensions: integration methods, condition states, and parameters sharing implementation
III-D Analysis of Multimodal Fusion Schemes
After defining the transformer backbone, the next step is to propose a fusion scheme that exploits the input reference image and audio to achieve the multimodality-guided video generation task. Given the condition modality, we summarize the existing fusion solutions into the following three schemes based on the transformer paradigm:
  • Direct Fusion directly feeds the condition modality into each block of the backbone and fuses it with the inner features via cross-attention;
  • Siamese Fusion feeds the condition modality into another similar transformer, extracts features from each block, and fuses them with the features in the corresponding block of the backbone;
  • Symbiotic Fusion concatenates condition modality with inputs before feeding it into the backbone and achieving seamless fusion via inherent self-attention.
Integration methods: The choice between self-attention and cross-attention involves trade-offs in computational efficiency and optimization. Cross-attention, with smaller interaction matrices, is computationally lighter and easier to optimize, especially in limited data scenarios, as demonstrated in early text-to-image models [13, 52]. In contrast, self-attention, which requires larger matrices for multimodal token alignment, is more challenging to optimize but achieves higher embedding quality with sufficient data [53]. Alternatively, addition, which sums hidden states, provides a simpler and more computationally efficient approach, particularly useful when conditional inputs aim to highlight specific local features in the latent space.
Condition States: The static versus dynamic nature of conditions reflects their constraint strength. Dynamic states, as in Siamese and Symbiotic Fusion, enable hierarchical refinement of conditional features during training, aligning with findings from ReferenceNet [17] and ControlNet [14] that multi-level conditioning enhances generation controllability. This flexibility improves feature alignment and model performance.
Parameter Sharing: Siamese Fusion uses separate network parameters for conditions and latent noise, while Symbiotic Fusion shares the same parameters. Direct Fusion, relying solely on cross-attention, also does not share parameters. Parameter sharing in Symbiotic Fusion is particularly advantageous for large-scale training, improving efficiency and scalability.
Modality Adaptation. The adaptability of fusion schemes to different modalities is key to their success in multimodal tasks. Direct Fusion, though offering weak alignment, is highly cost-effective for modalities with significant differences, such as text and image, making it ideal for resource-constrained scenarios. Siamese Fusion offers enhanced design flexibility by allowing adjustable modality alignment. This is exemplified in models such as ControlNet [14], which employs visual cue-level alignment, and AnimateAnyone [17], which utilizes pixel-level alignment. The strength of alignment can thus be precisely tailored to meet the specific requirements of the task at hand. Symbiotic Fusion enhances parameter efficiency compared to Siamese Fusion, achieving strong alignment with fewer parameters. This makes it particularly effective for large-scale applications, such as AnimateAnyone [17], where shared parameters ensure robust and efficient multimodal integration. The comparative analysis of the three fusion approaches is summarized in Table I.
In summary, we adopt Symbiotic Fusion for portrait features to ensure strong visual consistency, and Direct Fusion for audio features to maintain motion diversity and effective audio-visual alignment.
III-E Guidance on Reference Portrait and Audio
[…]
As for the audio, we require it to reasonably drive the animation, especially the lip motion [1]. However, certain unconscious actions (e.g. blinking and head movements) are almost independent of the audio yet serve to increase diversity. Considering that the modality of audio is quite different from the video, using Siamese Fusion and Symbiotic Fusion often causes ambiguity in the backbone learning, damaging the diversity. In a nutshell, Direct Fusion is concise yet effective for fusing the input audio.
For the audio representation, to accommodate the tokenization mechanism of the transformer, we follow [18] and extract richer semantic information from different layers of the wav2vec encoder [54]. Through nonlinear mapping, we obtain audio tokens $\bm{C}\in\mathbb{R}^{F\times T\times d^{\prime}}$ enriched with semantic content, which are aligned with latent features in the same feature space. Temporal-level audio representations are then synchronized with video sequences, resulting in the semantic audio representation. The experiments in Section IV-F verify the efficacy of our mechanism in feature fusion.
IV Experiment
Figure 5: The qualitative comparisons with other cutting-edge methods on the HDTF dataset. Our method achieves better audio-animation alignment (e.g. lip motions) and produces expressive results.
Figure 6: The qualitative comparisons with the existing portrait image animation approaches on the CelebV-HQ dataset. Our method achieves better portrait consistency.
IV-A Experimental Setups
Datasets We primarily conduct comprehensive experiments on two public datasets: HDTF [55] and CelebV-HQ [56]. These datasets include videos that consist of individuals of various ages, ethnicities, and genders, captured in half-body or close-up shots against varied indoor and outdoor backgrounds. To ensure high-quality training data, we adopt the preprocess scheme used in [18]. We first exclude videos with scene changes, significant camera movements, excessive facial motion, and side-facing shots. Then we refine the remaining videos and retain only single-person speaking videos with strong audio-lip correspondence. After refinement, we obtain 356 clips with 13.95 hours from the HDTF, and 24438 clips with 44.91 hours from the CelebV-HQ dataset. During training, we extract 16-frame video clips from these datasets using a specific sampling interval, each frame is resized to $512\times 512$ resolution.
Evaluation Metrics The evaluation metrics utilized in the portrait image animation approach include Fréchet Inception Distance (FID), Fréchet Video Distance (FVD), Structural Similarity Index Measure (SSIM) and Expression-FID (E-FID). Specifically, FID and FVD measure the similarity between generated images and real data, with lower values indicating better performance and thus more realistic outputs [57]. SSIM indices measure the structural similarity between ground truth videos and generated videos. E-FID leverages Inception features to assess image authenticity, providing a more precise fidelity measure. It extracts facial expression parameters via a reconstruction method [58] and computes FID to quantify the gap between generated and ground truth expressions.
Baseline In experiments, we compare our method with the publicly available implementations of DreamTalk (34M) [33], SadTalker (138M) [1], AniPortrait (2.8B) [2], Hallo (2.4B) [18], Hallo 2 (2.4B) [20] and EchoMimic (2.6B) [19]. The evaluation is performed on the HDTF and CelebV-HQ Dataset, utilizing a training and testing split where $90\%$ of the identity data was allocated for training purposes.
Figure 7: English audio-driven results of LetsTalk. Note that each column corresponds to the same audio.
Method FID $\downarrow$ FVD $\downarrow$ SSIM $\uparrow$ E-FID $\downarrow$
DreamTalk [33] 105.434 529.510 0.406 3.466
SadTalker [1] 102.371 362.634 0.414 2.168
AniPortrait [2] 78.284 331.117 0.623 2.903
EchoMimic [19] 55.791 299.343 0.639 2.493
Hallo [18] 45.754 285.988 0.698 2.235
Hallo2 [20] 44.343 241.628 0.657 2.017
Ours (Base) 46.240 289.385 0.666 2.587
Ours (Large) 31.761 219.119 0.632 1.962
TABLE II: The quantitative comparisons with the existed portrait image animation approaches on the HDTF dataset.
Method FID $\downarrow$ FVD $\downarrow$ SSIM $\uparrow$ E-FID $\downarrow$
DreamTalk [33] 64.297 293.905 0.448 5.915
SadTalker [1] 42.017 77.802 0.469 3.372
AniPortrait [2] 33.408 107.236 0.578 3.722
EchoMimic [19] 12.932 59.944 0.567 3.671
Hallo [18] 8.504 46.431 0.580 3.280
Hallo2 [20] 8.734 41.213 0.543 3.128
Ours (Base) 9.585 36.159 0.598 3.349
Ours (Large) 7.610 33.175 0.584 3.038
TABLE III: The quantitative comparisons with the existed portrait image animation approaches on the CelebV-HQ dataset.
IV-B Additional Implementation Details
Unlike previous methods [36, 18, 2] that use a two-stage training process, our approach adopts an end-to-end training strategy. We provide two model sizes: a Base model for simple fine-tuning, and a Large model with better generalization.
Diffusion We follow the hyperparameters from ADM [12], using a linear variance schedule with $t_{\text{max}}=1000$, ranging from $1\times 10^{-4}$ to $2\times 10^{-2}$. We also use ADM’s parameterization of the covariance $\Sigma_{\theta}$ and their method for timestep embedding.
Training LetsTalk is initialized from a pre-trained Latte [15] model. We use gradient clipping for stability and the AdamW optimizer with a fixed learning rate of $1\times 10^{-4}$ and no weight decay, following Latte’s training setup. As in previous generative modeling work [29, 59], we maintain an exponential moving average (EMA) of model weights with a decay rate of $0.9999$, and report all results using the EMA model.
Inference At inference, the model takes a single reference portrait and driving audio as input, and generates a video sequence that animates the reference image according to the audio. The generated video has the same length and resolution as during training: $16$ frames at $512\times 512$ resolution.
IV-C Quantitative Results
We report comprehensive quantitative comparisons on the HDTF and CelebV-HQ datasets in Table II and Table III, respectively. As shown in the tables, our method achieves the best performance on most metrics. Specifically, on the HDTF dataset, LetsTalk-L achieves the lowest FID (31.761) and FVD (219.119), outperforming the previous best method Hallo2 by 28.3% and 9.3%, respectively. On CelebV-HQ, our method also achieves the lowest FID (7.610) and FVD (33.175). These results demonstrate that our approach generates more realistic and temporally consistent talking-head videos. All methods generate the first video clip per validation sample and are evaluated at the same resolution.
IV-D Qualitative Results
We have visualized the results of different methods on the HDTF dataset (Fig. 5) and the CelebV-HQ dataset (Fig. 6). From the visualization results, it is evident that our method performs exceptionally well on both datasets. In comparison to the visual cues most strongly associated with audio (such as facial muscle and lip motions), our results closely align with the standard Ground Truth. Benefitting from our Symbiotic Fusion, our generated videos maintain strict portrait consistency. Furthermore, by employing Direct Fusion of audio, the characters exhibit strong expressiveness and diversity in our generated videos while achieving audio-animation alignment.
Figure 8: Chinese audio-driven results of LetsTalk. Note that each column corresponds to the same audio.
Method #Params Inference Time (ms/frame) FID $\downarrow$
SadTalker [1] 138M 23 93
EchoMimic [19] 2.2B 54 55
AniPortrait [2] 2.6B 74 82
Hallo [18] 2.4B 115 44
LetsTalk-B (Ours) 0.3B 27 46
LetsTalk-L (Ours) 0.7B 40 37
TABLE IV: Comparison of model size, inference time (ms per frame), and FID on the HDTF dataset.
IV-E Efficiency Comparison
To further demonstrate the practical value of our method, we compare model size, inference speed, and FID score with previous state-of-the-art methods. As shown in Table IV, LetsTalk-B achieves comparable or better visual quality with significantly fewer parameters and faster inference than recent diffusion-based approaches. The larger LetsTalk-L further improves FID while maintaining reasonable efficiency. These results highlight the strong trade-off between quality and efficiency achieved by LetsTalk, making it well-suited for real-world applications.
IV-F Ablation Study
In this section, we analyze the impact of different fusion schemes as illustrated in Fig. 4 on guidance using reference portrait and audio. Specifically, for analyzing Siamese Fusion, we have to design Siamese transformers for both reference portrait and audio to produce guidance for each block of the backbone. For the reference portrait, we leverage a DiT model pre-trained on ImageNet. For the audio, we define a transformer whose blocks are equipped with two-layer MLPs. The parameters of these transformers are simultaneously updated during training.
Reference Portrait FID $\downarrow$ FVD $\downarrow$ SSIM $\uparrow$ E-FID $\downarrow$
Direct Fusion 52.999 389.01 0.497 3.149
Siamese Fusion 47.779 286.692 0.651 2.597
Symbiotic Fusion 46.240 289.385 0.666 2.587
TABLE V: Experimental results of three multimodal fusion schemes, guided by reference portraits, on the HDTF validation dataset, employing the Letstalk-B model.
Driven Audio FID $\downarrow$ FVD $\downarrow$ SSIM $\uparrow$ E-FID $\downarrow$
Symbiotic Fusion 51.880 403.628 0.471 3.398
Siamese Fusion 48.041 318.637 0.585 2.706
Direct Fusion 46.240 289.385 0.666 2.587
TABLE VI: Experimental results of three multimodal fusion schemes, driven by audio inputs, on the HDTF validation dataset, employing the Letstalk-B model.
Analysis on the Reference Portrait Fusion Portrait consistency is crucial to the portrait image animation task and highly relevant to the portrait fusion scheme used. We evaluate the synthesis quality of portrait fusion using three schemes, and the quantitative results are shown in Table V. Symbiotic Fusion almost surpasses the other two schemes, realizing better portrait consistency and video quality. Besides, while Siamese Fusion produces comparable results, it introduces additional parameters (i.e. siamese transformer) and complicates the learning task.
Analysis on the Driven Audio Fusion The way we fuse audio affects both audio-animation alignment and diversity. Audio-animation alignment refers to specific visual cues [1] (e.g. facial muscle and lip motions), whereas the diversity is related to the speech-agnostic motions (e.g. blinking and neck moving) that often appear as unconscious actions. We observe that diversity mainly comes from the transformer backbone during denoising. Thus, we aim to fuse the audio to achieve alignment while preserving the backbone’s inherent diversity. Quantitative results in Table VI indicate that Direct Fusion outperforms the other schemes in establishing the correspondence between audio and animation.
Figure 9: Portrait animation results without audio guidance for three fusion schemes. Siamese and Symbiotic fusion schemes with high fusion compactness limit the diversity of video generation, while Direct fusion preserves this diversity.
Impact of Audio-guidance on Diversity To validate the impact of different fusion schemes on diversity, we serve LetsTalk as an unconditional video generator [^1] Filling audio embeddings with 0 for fusion. and compare the diversity of character actions in the generated videos. The results in Fig. 9 demonstrate that the deeper the fusion compactness, the more static the generated video. The tightest Symbiotic Fusion generates a portrait with almost no motions, Siamese Fusion produces slight motions, while the least tight Direct Fusion we adopt results in natural motions.
IV-G User Study
To comprehensively assess the perceptual quality of our approach, we conducted a user study with 12 participants. Each participant was shown videos generated by three different methods: Hallo, EchoMimic, and our proposed LetsTalk-L. For the evaluation, 20 audio samples were randomly selected, and for each audio, participants were asked to rank the three corresponding videos according to two criteria: realism, which measures how natural and lifelike the video appears, and synchronization, which evaluates how well the lip movements match the audio.
The results of the user study are summarized in Fig 10. The table reports the percentage of times each method was ranked first, second, or third for both realism and synchronization. As shown in the results, LetsTalk-L achieves the highest proportion of first-place rankings in both categories. These findings indicate that users consistently preferred the videos generated by our method, both in terms of visual realism and the accuracy of lip synchronization with the audio. Overall, the user study demonstrates the effectiveness of our approach in producing high-quality, realistic, and well-synchronized talking head videos.
Figure 10: User study results (%) on (a) realism (left) and (b) synchronization (right).
Figure 11: Singing audio-driven results of LetsTalk. Note that each column corresponds to the same audio.
IV-H Long-Duration Generation
To validate the effectiveness of our proposed approach for long video generation, we conducted experiments to analyze the quality of 100 consecutive video clips (about 64 second) generated under three settings: (1) the baseline model (LetsTalk-L) with with the memory bank, (2) the method augmented with regularized training strategy, and (3) the method further enhanced with the patch drop augmentation. Our results in Table VII demonstrate that the memory bank significantly improves temporal consistency, while the regularized training strategy further mitigates degradation in later clips, leading to higher overall video quality. The results demonstrate that regularized training strategy primarily enhances single-frame quality, significantly reducing FID from 52.243 to 32.746, while maintaining temporal consistency (FVD 264.709). On the other hand, patch drop augmentation further improves long-term motion coherence, achieving the best FVD (251.684) with only a slight FID increase (34.821), confirming its complementary role in stabilizing long-duration generation.
[…]
52.243 276.16
43.728 263.173
32.746 264.709
34.821 251.684
TABLE VII: Experimental results of long video generation. Experimental results comparing three approaches for 64-second video generation (100 clips): (1) memory bank baseline, (2) regularized training, and (3) patch drop augmentation. ✓ indicates the component is enabled.
IV-I Visualization
We provide video visualizations showcasing long-duration results of up to 2 minutes. In these videos, each row corresponds to a different portrait driven by the same audio segment, demonstrating our model’s ability to generate consistent and diverse talking head videos across various identities. Notably, our model is trained only on HDTF [55] and CelebV-HQ [56], which feature English conversational scenarios with real portraits. To further demonstrate the generalization ability of our approach, we also present results where real portraits are driven by Chinese speech and singing audio, even though such scenarios are not included in the training data.
V Conclusion
In this work, we present LetsTalk, a novel diffusion transformer framework for efficient and high-quality long-duration talking video synthesis under multimodal guidance. LetsTalk introduces a memory bank mechanism and noise-regularized training to explicitly maintain contextual continuity and mitigate error accumulation in extended video generation. Our spatiotemporal modeling, based on a deep compression autoencoder and linear attention, achieves state-of-the-art visual fidelity with substantially fewer parameters and faster inference. Through systematic analysis of multimodal fusion, we provide principled guidance for multimodal conditioning in video generation, demonstrating that deep Symbiotic Fusion for portrait features and shallow Direct Fusion for audio ensures both strong identity preservation and expressive, synchronized motion. Extensive experiments show that LetsTalk consistently outperforms existing approaches in quality, temporal coherence, and efficiency, while generalizing robustly to out-of-domain scenarios. Overall, LetsTalk advances the frontier of realistic, scalable talking video generation, paving the way for broader applications in digital humans and interactive media.