### v4

# Efficient Long-duration Talking Video Synthesis with Linear Diffusion Transformer under Multimodal Guidance

Abstract
Long-duration talking video synthesis faces enduring challenges in achieving high video quality, portrait and temporal consistency, and computational efficiency. As video length increases, issues such as visual degradation, identity inconsistency, temporal incoherence, and error accumulation become increasingly problematic, severely affecting the realism and reliability of the results. To address these challenges, we present LetsTalk, a diffusion transformer framework equipped with multimodal guidance and a novel memory bank mechanism, explicitly maintaining contextual continuity and enabling robust, high-quality, and efficient generation of long-duration talking videos. In particular, LetsTalk introduces a noise-regularized memory bank to alleviate error accumulation and sampling artifacts during extended video generation. To further improve efficiency and spatiotemporal consistency, LetsTalk employs a deep compression autoencoder and a spatiotemporal-aware transformer with linear attention for effective multimodal fusion. We systematically analyze three fusion schemes and show that combining deep (Symbiotic Fusion) for portrait features and shallow (Direct Fusion) for audio achieves superior visual realism and precise speech-driven motion, while preserving diversity of movements. Extensive experiments demonstrate that LetsTalk establishes new state-of-the-art in generation quality, producing temporally coherent and realistic talking videos with enhanced diversity and liveliness, and maintains remarkable efficiency with 8 fewer parameters than previous approaches.
I Introduction
Animating a portrait image with speech, commonly referred to as talking head generation, aims to synthesize a speaking portrait from a single image and accompanying audio. This technology has promising applications in gaming, film production, and digital marketing. Despite recent advancements, long-duration talking video synthesis continues to grapple with maintaining high video quality, portrait consistency, temporal coherence, and computational efficiency [1, 2, 3, 4]. As video duration increases, issues such as visual degradation, loss of identity, and error accumulation become more pronounced, diminishing overall realism and reliability.
Traditional approaches mainly rely on 3D morphable models (3DMM) [5, 6, 7], which deliver temporal and identity consistency but are often limited in their expressiveness [1, 2]. Audio-to-2D facial motion techniques [8, 9, 10, 11, 12], along with multimodal guidance strategies such as Anyonenet [3] and TalkCLIP [4], enrich animation diversity and introduce control via text or voice. Recent progress further enables personalized head movement [13] and text-audio-guided video generation [14], but existing methods often struggle to effectively disentangle static and dynamic factors, thereby sometimes compromising consistency or limiting expressiveness across longer time spans.
Diffusion models [15, 16, 17] have recently achieved remarkable success in generating realistic 2D images [17, 18] and videos [19, 20]. AnimateAnyone [21] showcases the strength of motion modules and ReferenceNet for pose-controllable human animation within diffusion frameworks. Building on these advances, recent works [22, 23, 24] directly employ audio as guidance to synthesize talking head videos, leveraging diffusion models to improve both diversity and expressiveness.
Nevertheless, significant challenges remain for long-duration talking head generation. As video length increases, existing methods frequently struggle to preserve visual quality, portrait consistency, and temporal coherence. Accumulated noise and artifacts, such as distortions and flickering, further undermine realism. Efficiency is often constrained by reliance on auxiliary spatial signals for audio control, unnecessary parameter overhead from fine-tuning generic models for animation tasks, and computationally intensive architectures, including low-compression autoencoders and standard attention schemes. Additionally, effective strategies for fusing multimodal conditions, such as reference portraits and driving audio, have not been sufficiently explored in conditional video generation frameworks.
In this work, we introduce LetsTalk, a diffusion transformer framework crafted for efficient, high-quality generation of realistic, long-duration talking videos under multimodal conditions. A central innovation in LetsTalk for long video generation is the Noise-Regularized Memory Bank, which enables each generated clip to access information from prior clips while explicitly simulating inference-time noise accumulation during training. This mechanism effectively preserves contextual continuity and mitigates error accumulation and sampling artifacts throughout extended video sequences, ensuring smooth transitions and high temporal consistency. To further improve efficiency and scalability, LetsTalk incorporates a Deep Compression Autoencoder that compresses input frames by 32 in the latent space, drastically reducing the number of tokens while retaining essential visual features. Transformer blocks are equipped with specialized temporal and spatial modules, utilizing Linear Attention [25] for efficient modeling of both intra-frame and inter-frame coherence without excessive computational cost. This design empowers LetsTalk to sustain high generation quality and temporal consistency over long video durations.
For multimodal guidance, defined here as fusing reference portraits and driving audio for video generation, we first categorize three fusion schemes by their integration tightness: 1) Direct Fusion, 2) Siamese Fusion, and 3) Symbiotic Fusion. These schemes are analyzed in terms of integration depth, conditional state, and parameter sharing, providing a systematic summary of their adaptability to diverse modalities. We further examine the underlying factors that influence such adaptability, addressing theoretical gaps in conditional guidance modeling. For effective long-duration video generation, we employ Symbiotic Fusion with the reference image to ensure strong portrait consistency across extended sequences due to its visual proximity to the generated video, and use Direct Fusion for audio to tightly synchronize lip movement and expressions with speech while fostering diversity in other movements (e.g., blinking, head turns) across longer durations.
We summarize the contributions of this work as follows:
We present LetsTalk, a novel and efficient diffusion transformer framework for scalable, long-duration talking video generation, featuring a powerful spatiotemporal modeling architecture that achieves high video quality and computational efficiency with significantly fewer parameters.
We propose a noise-regularized memory bank mechanism that preserves contextual continuity and effectively curbs error accumulation, substantially improving temporal consistency in extended video sequences.
We provide a systematic analysis of multimodal fusion schemes and demonstrate that combining deep Symbiotic Fusion for portrait features and Direct Fusion for audio achieves superior visual quality and temporal coherence.
Experiments demonstrate that our method achieves realistic, coherent, and diverse long-duration talking videos with optimal quality-efficiency trade-offs.
II Related Work
II-A Diffusion Models for Video Generation
Diffusion models have demonstrated exceptional potential across various domains, particularly in image [15, 16] and video generation [26, 21, 27, 28]. A notable example is Stable Diffusion [17], which utilizes a UNet architecture to iteratively produce high-quality images from text prompts after extensive training on large-scale text-image datasets [29]. This approach has been extended to more complex tasks like video generation [26, 21], where models such as Video Diffusion Models (VDM)[27] apply the diffusion process with space-time factorized training, and ImagenVideo[30] enhances VDM with cascaded diffusion models for high-resolution outputs. Further advances, including VideoComposer [31], AnimateDiff [26], and VideoCrafter [32], focus on synthesizing videos by conditioning the diffusion process on images and blending image latents. A significant direction in the evolution of diffusion models is the incorporation of patch sequences of images and transformers, exemplified by the Diffusion Transformer (DiT) architecture [33], which facilitates scaling to larger datasets and model parameters. Recent works, such as Latte [19] and VDT [20], have extended the DiT architecture for video generation by incorporating temporal mechanisms, achieving strong results by training diffusion models from scratch. Building on this, our study also adopts an end-to-end approach for high-fidelity, temporally consistent video generation.
II-B Audio-driven Portrait Image Animation
In recent years, the field of portrait image animation has advanced significantly, especially in generating realistic and expressive talking head animations from static images paired with audio inputs. Early efforts primarily focused on improving lip synchronization and frame generation [9, 8, 10]. The integration of 3D information and more expressive head movements in models such as SadTalker [1] and Genface [34] further improved realism. A major breakthrough comes from the adoption of diffusion models, exemplified by works like Diffused Heads [35], DreamTalk [36], and VividTalk [37], which produce high-fidelity talking head videos with enhanced lip synchronization and diverse expressions. Recent innovations have explored multimodal guidance and personalized control mechanisms. Notably, Anyonenet [3] pioneered synchronized speech and head generation for arbitrary identities, enabling voice and video synthesis from a single image. TalkCLIP [4] introduced text-guided expressive speaking styles, allowing users to control facial expressions via natural language. Yi et al. [13] proposed personalized head movement prediction, learning individual motion patterns from short video clips for more natural animation. TA2V [14] further explored text-audio guided video generation, leveraging multi-modal inputs for enhanced synthesis. Very recent works such as VASA-1 [38], EMO [39], and Hallo [22, 24] focus on emotionally expressive talking faces and advanced control over lip movements, expressions, and poses, pushing the boundaries of realism and user-controllability. These approaches, often employing latent diffusion frameworks pioneered in AnimateAnyone [21], demonstrate the rapid progress and increasing capabilities of audio-driven portrait image animation.
II-C Conditionally Guided Diffusion Model
The impressive generative capabilities of diffusion models have made the integration of control information during generation a crucial area of research. Text-guided control methods focus on refining prompts, manipulating CLIP features, and modifying cross-attention mechanisms [40, 41, 42, 43, 44]. ControlNet [18] enhances large pretrained text-to-image diffusion models by incorporating spatially localized, task-specific conditioning. In video generation, AnimateAnyone [21] introduces the ReferenceNet to ensure the consistency of fine-grained appearance details from reference images. Gen-1 [45] integrates structural guidance through depth maps and cross-attention, while MCDiff [46], LaMD [47], and MotionCtrl [48] focus on motion-guided video generation, enhancing the realism of human actions. For talking video synthesis, previous approaches primarily relied on image diffusion models to generate dynamic portraits frame by frame, starting from reference portraits [49, 35]. Recent methods often incorporate ReferenceNet for processing reference portraits and cross-attention mechanisms for handling audio information [39, 22]. While control over both images and videos in latent diffusion models has seen significant advancements, the theoretical understanding of various conditioning strategies remains limited, particularly in handling multiple modalities (e.g., image, video, audio). To address this gap, we investigate conditioning strategies in diffusion transformers, aiming to improve control and generation quality across modalities.
III Methodology
We propose LetsTalk, a diffusion transformer framework for audio-driven portrait animation, as depicted in Fig. 2. Section III-A introduces the efficient spatiotemporal transformer backbone. Section III-B details our mechanism for long-range consistency, including the memory bank and noise-regularized training. Section III-C analyzes multimodal fusion schemes, and Section III-D elaborates our multimodal guidance strategy.
III-A Efficient SpatioTemporal Transformer with Linear Attention
Deep Compression Autoencoder. We employ a deep compression autoencoder [25] that downsamples images by 32, achieving reconstruction quality comparable to SDXL while greatly reducing memory and computation costs. Separation of compression and denoising enables preservation of essential temporal cues, such as lip-sync dynamics. Compressed latent features are partitioned into non-overlapping patches and assigned spatial-temporal positional embeddings [50] to enforce structural consistency. This approach facilitates efficient transformer processing and further decreases memory usage by about four times compared to conventional methods.
Interleaved Attention Mechanism. Instead of conventional frame-by-frame processing as in DiT [33], we process latent features with interleaved spatial and temporal attention. Full-token attention incurs quadratic complexity , so we alternate spatial linear attention [51] and temporal softmax attention, reducing total cost to . Specifically, for spatial modeling, we reshape and apply ReLU-based linear attention:
where . For temporal modeling, features are transposed to , then subjected to softmax attention across the temporal dimension. This alternating strategy maintains global context and ensures scalability for high-resolution, long-duration inputs.
Mix-FFN Block. Although linear attention achieves similar final performance, its convergence may be slower. Following SANA [25], we replace the standard MLP-FFN with Mix-FFN, which integrates an inverted residual block, a depth-wise convolution, and a gated linear unit (GLU). This convolution enhances localization ability, compensating for the weaker locality observed with ReLU-based linear attention.
III-B Long-duration Generation with Noise-Regularized Memory Bank
In diffusion-based video generation, maintaining temporal alignment and consistency across long sequences is essential. However, as the video extends, the accumulation of residual noise and sampling artifacts can threaten temporal coherence and visual quality. This is further exacerbated by the mismatch between training and inference: during training, the memory bank is constructed from ground-truth frames, while during inference, it is updated with generated frames.
Memory Bank: To address these challenges, we introduce a memory bank that stores information from previously generated clips. Unlike previous frame-by-frame generation approaches [49, 35], our end-to-end diffusion transformer generates an entire video clip in a single sampling process. The memory bank is patchified and temporally transposed, resulting in , where denotes the number of frames selected from prior clips. The memory bank provides prior information to the current generation through Symbiotic Fusion with , leveraging the scalability of the Transformer architecture. This design ensures smooth transitions and temporal consistency across video segments, and is especially advantageous for long-duration video generation.
Theoretical Motivation: We observed that DDPM-based sampling cannot fully remove noise due to two main factors:
Model prediction error: The predicted noise only approximates the true noise , leaving some residual noise after denoising.
Sampling stochasticity: At each reverse step, new Gaussian noise is added:
(2) The term introduces inherent randomness.
Consequently, even after finite-step denoising, diffusion models retain residual noise, which accumulates as the video sequence extends. This motivates our noise-regularized memory bank, which differs from previous long video generation methods.
Noise-Regularized Training Strategy. To bridge the gap between training and inference, we propose a noise-regularized training strategy. We perturb the memory bank with controlled Gaussian noise, , simulating inference-time noise accumulation. Specifically, we randomly sample a diffusion timestep from the first 50 steps and add the corresponding noise to the memory bank during training. Additionally, inspired by Hallo2 [24], we apply patch-drop augmentation by randomly masking 25% of spatial patches in the memory bank, further enhancing model robustness.
This deliberate corruption enables the model to learn to correct for artifacts and maintain temporal dependencies. Our framework thus achieves robust and stable long-duration video synthesis, as demonstrated in Fig. 3.
| Schemes | Integration | State | Params | Modality Adapt. |
| Direct | Cross-attention | Static | Dedicated | Weak alignment |
| Siamese | Self-attention/Add | Dynamic | Dedicated | Flexible |
| Symbiotic | Self-attention | Dynamic | Shared | Strong |
III-C Analysis of Multimodal Fusion Schemes
After defining the transformer backbone, the next step is to propose a fusion scheme that exploits the input reference image and audio to achieve the multimodality-guided video generation task. Given the condition modality, we summarize the existing fusion solutions into the following three schemes based on the transformer paradigm:
Direct Fusion directly feeds the condition modality into each block of the backbone and fuses it with the inner features via cross-attention;
Siamese Fusion feeds the condition modality into another similar transformer, extracts features from each block, and fuses them with the features in the corresponding block of the backbone;
Symbiotic Fusion concatenates condition modality with inputs before feeding it into the backbone and achieving seamless fusion via inherent self-attention.
Integration methods: The choice between self-attention and cross-attention involves trade-offs in computational efficiency and optimization. Cross-attention, with smaller interaction matrices, is computationally lighter and easier to optimize, especially in limited data scenarios, as demonstrated in early text-to-image models [17, 52]. In contrast, self-attention, which requires larger matrices for multimodal token alignment, is more challenging to optimize but achieves higher embedding quality with sufficient data [53]. Alternatively, addition, which sums hidden states, provides a simpler and more computationally efficient approach, particularly useful when conditional inputs aim to highlight specific local features in the latent space.
Condition States: The static versus dynamic nature of conditions reflects their constraint strength. Dynamic states, as in Siamese and Symbiotic Fusion, enable hierarchical refinement of conditional features during training, aligning with findings from ReferenceNet [21] and ControlNet [18] that multi-level conditioning enhances generation controllability. This flexibility improves feature alignment and model performance.
Parameter Sharing: Siamese Fusion uses separate network parameters for conditions and latent noise, while Symbiotic Fusion shares the same parameters. Direct Fusion, relying solely on cross-attention, also does not share parameters. Parameter sharing in Symbiotic Fusion is particularly advantageous for large-scale training, improving efficiency and scalability.
Modality Adaptation. Fusion schemes must adapt to different modalities for effective multimodal tasks. Direct Fusion offers weak alignment but is highly efficient, making it suitable for modalities with large differences, such as text and image, especially in resource-limited settings. Siamese Fusion provides greater flexibility by allowing tunable alignment strength, as seen in ControlNet [18] (cue-level) and AnimateAnyone [21] (pixel-level), enabling precise control based on task demands. Symbiotic Fusion further improves parameter efficiency over Siamese Fusion, achieving strong alignment with fewer parameters—beneficial for large-scale applications like AnimateAnyone [21]. A comparison of these fusion approaches is shown in Table I.
In summary, we adopt Symbiotic Fusion for portrait features to ensure strong visual consistency, and Direct Fusion for audio features to maintain motion diversity and effective audio-visual alignment.
III-D Guidance on Reference Portrait and Audio
Given the reference image, we require that the character of the generated video is consistent with the reference. Since we adopt Symbiotic Fusion, leveraging the similarity between image and video modalities. The reference image is encoded into a portrait feature , temporally repeated across frames, and concatenated with video latent features to form , where . Shared self-attention in this fused space ensures consistent character representation across frames. In comparison, Direct Fusion leads to misalignment in deep layers, while Siamese Fusion adds complexity without significant benefits over Symbiotic Fusion.
For audio-driven animation, especially for lip motion [1], the modality gap between audio and video motivates the use of Direct Fusion, which is simple and preserves output diversity (e.g., blinking, head movements). Following [22], we extract multi-layer wav2vec [54] features and apply nonlinear mapping to obtain audio tokens enriched with semantic information. These are temporally synchronized with video features to guide animation. Ablation studies in Section IV-F validate the effectiveness of this fusion mechanism.
IV Experiment
IV-A Experimental Setups
Datasets We primarily conduct comprehensive experiments on two public datasets: HDTF [55] and CelebV-HQ [56]. These datasets include videos that consist of individuals of various ages, ethnicities, and genders, captured in half-body or close-up shots against varied indoor and outdoor backgrounds. To ensure high-quality training data, we adopt the preprocess scheme used in [22]. We first exclude videos with scene changes, significant camera movements, excessive facial motion, and side-facing shots. Then we refine the remaining videos and retain only single-person speaking videos with strong audio-lip correspondence. After refinement, we obtain 356 clips with 13.95 hours from the HDTF, and 24438 clips with 44.91 hours from the CelebV-HQ dataset. During training, we extract 16-frame video clips from these datasets using a specific sampling interval, each frame is resized to resolution.
Evaluation Metrics The evaluation metrics utilized in the portrait image animation approach include Fréchet Inception Distance (FID), Fréchet Video Distance (FVD), Structural Similarity Index Measure (SSIM) and Expression-FID (E-FID). Specifically, FID and FVD measure the similarity between generated images and real data, with lower values indicating better performance and thus more realistic outputs [57]. SSIM indices measure the structural similarity between ground truth videos and generated videos. E-FID leverages Inception features to assess image authenticity, providing a more precise fidelity measure. It extracts facial expression parameters via a reconstruction method [58] and computes FID to quantify the gap between generated and ground truth expressions.
Baseline In experiments, we compare our method with the publicly available implementations of DreamTalk (34M) [36], SadTalker (138M) [1], AniPortrait (2.8B) [2], Hallo (2.4B) [22], Hallo 2 (2.4B) [24] and EchoMimic (2.6B) [23]. The evaluation is performed on the HDTF and CelebV-HQ Dataset, utilizing a training and testing split where of the identity data was allocated for training purposes.
| Method | FID | FVD | SSIM | E-FID |
| DreamTalk [36] | 105.434 | 529.510 | 0.406 | 3.466 |
| SadTalker [1] | 102.371 | 362.634 | 0.414 | 2.168 |
| AniPortrait [2] | 78.284 | 331.117 | 0.623 | 2.903 |
| EchoMimic [23] | 55.791 | 299.343 | 0.639 | 2.493 |
| Hallo [22] | 45.754 | 285.988 | 0.698 | 2.235 |
| Hallo2 [24] | 44.343 | 241.628 | 0.657 | 2.017 |
| Ours (Base) | 46.240 | 289.385 | 0.666 | 2.587 |
| Ours (Large) | 31.761 | 219.119 | 0.632 | 1.962 |
| Method | FID | FVD | SSIM | E-FID |
| DreamTalk [36] | 64.297 | 293.905 | 0.448 | 5.915 |
| SadTalker [1] | 42.017 | 77.802 | 0.469 | 3.372 |
| AniPortrait [2] | 33.408 | 107.236 | 0.578 | 3.722 |
| EchoMimic [23] | 12.932 | 59.944 | 0.567 | 3.671 |
| Hallo [22] | 8.504 | 46.431 | 0.580 | 3.280 |
| Hallo2 [24] | 8.734 | 41.213 | 0.543 | 3.128 |
| Ours (Base) | 9.585 | 36.159 | 0.598 | 3.349 |
| Ours (Large) | 7.610 | 33.175 | 0.584 | 3.038 |
IV-B Additional Implementation Details
Unlike previous methods [39, 22, 2] that use a two-stage training process, our approach adopts an end-to-end training strategy. We provide two model sizes: a Base model for simple fine-tuning, and a Large model with better generalization.
Diffusion We follow the hyperparameters from ADM [16], using a linear variance schedule with , ranging from to . We also use ADM’s parameterization of the covariance and their method for timestep embedding.
Training LetsTalk is initialized from a pre-trained Latte [19] model. We use gradient clipping for stability and the AdamW optimizer with a fixed learning rate of and no weight decay, following Latte’s training setup. As in previous generative modeling work [33, 59], we maintain an exponential moving average (EMA) of model weights with a decay rate of , and report all results using the EMA model.
Inference At inference, the model takes a single reference portrait and driving audio as input, and generates a video sequence that animates the reference image according to the audio. The generated video has the same length and resolution as during training: frames at resolution.
IV-C Quantitative Results
We report comprehensive quantitative comparisons on the HDTF and CelebV-HQ datasets in Table II and Table III, respectively. As shown in the tables, our method achieves the best performance on most metrics. Specifically, on the HDTF dataset, LetsTalk-L achieves the lowest FID (31.761) and FVD (219.119), outperforming the previous best method Hallo2 by 28.3% and 9.3%, respectively. On CelebV-HQ, our method also achieves the lowest FID (7.610) and FVD (33.175). These results demonstrate that our approach generates more realistic and temporally consistent talking-head videos. All methods generate the first video clip per validation sample and are evaluated at the same resolution.
IV-D Qualitative Results
We have visualized the results of different methods on the HDTF dataset (Fig. 5) and the CelebV-HQ dataset (Fig. 6). From the visualization results, it is evident that our method performs exceptionally well on both datasets. In comparison to the visual cues most strongly associated with audio (such as facial muscle and lip motions), our results closely align with the standard Ground Truth. Benefitting from our Symbiotic Fusion, our generated videos maintain strict portrait consistency. Furthermore, by employing Direct Fusion of audio, the characters exhibit strong expressiveness and diversity in our generated videos while achieving audio-animation alignment.
IV-E Efficiency Comparison
To further demonstrate the practical value of our method, we compare model size, inference speed, and FID score with previous state-of-the-art methods. As shown in Table IV, LetsTalk-B achieves comparable or better visual quality with significantly fewer parameters and faster inference than recent diffusion-based approaches. The larger LetsTalk-L further improves FID while maintaining reasonable efficiency. These results highlight the strong trade-off between quality and efficiency achieved by LetsTalk, making it well-suited for real-world applications.
IV-F Ablation Study
In this section, we analyze the impact of different fusion schemes as illustrated in Fig. 4 on guidance using reference portrait and audio.
| Reference Portrait | FID | FVD | SSIM | E-FID |
| Direct Fusion | 52.999 | 389.01 | 0.497 | 3.149 |
| Siamese Fusion | 47.779 | 286.692 | 0.651 | 2.597 |
| Symbiotic Fusion | 46.240 | 289.385 | 0.666 | 2.587 |
| Driven Audio | FID | FVD | SSIM | E-FID |
| Symbiotic Fusion | 51.880 | 403.628 | 0.471 | 3.398 |
| Siamese Fusion | 48.041 | 318.637 | 0.585 | 2.706 |
| Direct Fusion | 46.240 | 289.385 | 0.666 | 2.587 |
Analysis on the Reference Portrait Fusion Portrait consistency is crucial to the portrait image animation task and highly relevant to the portrait fusion scheme used. We evaluate the synthesis quality of portrait fusion using three schemes, and the quantitative results are shown in Table V. Symbiotic Fusion almost surpasses the other two schemes, realizing better portrait consistency and video quality. Besides, while Siamese Fusion produces comparable results, it introduces additional parameters (i.e. siamese transformer) and complicates the learning task.
Analysis on the Driven Audio Fusion The way we fuse audio affects both audio-animation alignment and diversity. Audio-animation alignment refers to specific visual cues [1] (e.g. facial muscle and lip motions), whereas the diversity is related to the speech-agnostic motions (e.g. blinking and neck moving) that often appear as unconscious actions. We observe that diversity mainly comes from the transformer backbone during denoising. Thus, we aim to fuse the audio to achieve alignment while preserving the backbone’s inherent diversity. Quantitative results in Table VI indicate that Direct Fusion outperforms the other schemes in establishing the correspondence between audio and animation.
Impact of Audio-guidance on Diversity To validate the impact of different fusion schemes on diversity, we serve LetsTalk as an unconditional video generator111Filling audio embeddings with for fusion. and compare the diversity of character actions in the generated videos. The results in Fig. 7 demonstrate that the deeper the fusion compactness, the more static the generated video. The tightest Symbiotic Fusion generates a portrait with almost no motions, Siamese Fusion produces slight motions, while the least tight Direct Fusion we adopt results in natural motions.
IV-G User Study
We conducted a user study with 12 participants to assess the perceptual quality of our approach. Each participant viewed videos generated by Hallo, EchoMimic, and our proposed LetsTalk-L, using 20 randomly selected audio samples. For each sample, participants ranked the three methods by realism (naturalness) and synchronization (lip-audio match). The results of the user study are summarized in Fig 8. The table reports the percentage of times each method was ranked first, second, or third for both realism and synchronization. As shown in the results, LetsTalk-L achieves the highest proportion of first-place rankings in both categories. LetsTalk-L consistently received the most first-place votes in both categories, indicating clear user preference. This demonstrates that our method produces videos with superior visual realism and lip synchronization.
IV-H Long-Duration Generation
We evaluated our approach for long video generation using 100 consecutive clips ( 64 seconds) under three settings: (1) baseline LetsTalk-L with memory bank, (2) regularized training, and (3) additional patch drop augmentation. As shown in Table VII, the memory bank improves temporal consistency, regularized training boosts single-frame quality (FID drops from 52.243 to 32.746, FVD 264.709), and patch drop augmentation further enhances motion coherence, achieving the lowest FVD (251.684) with minimal FID increase (34.821). These results confirm that each technique effectively improves long-duration video quality.
| Memory Bank | Patch Drop | Regularized Training | FID | FVD |
| ✓ | 52.243 | 276.16 | ||
| ✓ | ✓ | 43.728 | 263.173 | |
| ✓ | ✓ | 32.746 | 264.709 | |
| ✓ | ✓ | ✓ | 34.821 | 251.684 |
V Conclusion
In this work, we present LetsTalk, a novel diffusion transformer framework for efficient and high-quality long-duration talking video synthesis under multimodal guidance. LetsTalk introduces a noise-regularized memory bank to explicitly maintain contextual continuity and mitigate error accumulation in extended video generation. Our spatiotemporal modeling, based on a deep compression autoencoder and linear attention, achieves state-of-the-art visual fidelity with substantially fewer parameters and faster inference. Through systematic analysis of multimodal fusion, we provide principled guidance for multimodal conditioning in video generation, demonstrating that deep Symbiotic Fusion for portrait features and shallow Direct Fusion for audio ensures both strong identity preservation and expressive, synchronized motion. Extensive experiments show that LetsTalk consistently outperforms existing approaches in quality, temporal coherence, and efficiency, while generalizing robustly to out-of-domain scenarios. Overall, LetsTalk advances the frontier of realistic, scalable talking video generation, paving the way for broader applications in digital humans and interactive media.